### v3

# Central limit theorems for interacting innovation processes, related statistical tools and general results

Abstract.
We study a networked system of innovation processes, where each process is modeled as an urn with infinitely many colors-a classical framework for capturing the emergence of novelties. Extending this paradigm, we analyze a model of interacting urns, where the probability of generating or reusing elements in one process is influenced by the histories of others. This interaction is governed by two matrices that control innovation triggering and reinforcement dynamics across the system. The core contribution of this work is a detailed analysis of the second-order asymptotic behavior of the model. Building on these theoretical results, we develop statistical tools to infer the structure and strength of inter-process influence. The methodology is framed in a general setting, making it broadly applicable. We validate our approach with applications to two real-world datasets from Reddit discussions and Gutenberg text corpora.
Key-words: Reinforcement, Interaction, Urn model, Poisson-Dirichlet process, Innovation process, Central limit theorem, stable convergence.
G. Aletti is a member of “Gruppo Nazionale per il Calcolo Scientifico (GNCS)” of the Italian Institute “Istituto Nazionale di Alta Matematica (INdAM)”. I. Crimaldi and A. Ghiglietti are members of “Gruppo Nazionale per l’Analisi Matematica, la Probabilità e le loro Applicazioni (GNAMPA)” of the Italian Institute “Istituto Nazionale di Alta Matematica (INdAM)”. Irene Crimaldi thanks the “Resilienza Economica e Digitale” (RED) project (CUP D67G23000060001) funded by the Italian Ministry of University and Research (MUR) as “Department of Excellence” (Dipartimenti di Eccellenza 2023-2027, Ministerial Decree no. 230/2022). Partially supported by the European Union - NextGenerationEU through the Italian Ministry of University and Research under the National Recovery and Resilience Plan (PNRR) - M4C2 - Investment 1.3, title [Public Sector Indicators for Sustainability and Wellbeing (PUWELL)] - Program [Growing Resilient, INclusive and Sustainable (GRINS)] - PE18 - CUP J33C22002910001.
Understanding the mechanisms by which novelties emerge, propagate, and catalyze further innovations is pivotal across various disciplines, including biology, linguistics, and social sciences Armano and Javarone (2017); Arthur (2009); Fink et al. (2017); Puglisi et al. (2008); Rogers et al. (2019); Rzhetsky et al. (2015)). In probabilistic terms, a novelty or innovation is defined as the first occurrence of an event of interest.
A widely adopted mathematical model for such innovation process is an urn model with infinitely many colors Bandyopadhyay and Thacker (2022); Sariev et al. (2023), also known as species sampling sequence Favaro et al. (2009); Hansen and Pitman (2000); Lijoi et al. (2007); Lijoi and Prünster (2010); Pitman (1996). For each $t\in{\mathbb{N}}\setminus\{0\}$, let $C_{t}$ denote the observed color at extraction $t$. The probability that the $(t+1)$-th draw yields a “new” color (i.e., one not previously observed) is denoted by $Z^{*}_{t}$ and it is a function of the observations $C_{1},\dots,C_{t}$. Conversely, the probability that the next color is an “old” (i.e., already observed) color $c$ is given by $P_{c,t}=\sum_{n=1}^{t}Q_{n,t}\mathbb{I}_{\{C_{n}=c\}}$, where $Q_{n,t}$ depends on the observations $C_{1},\dots,C_{t}$. The quantities $Z^{*}_{t}$ and $Q_{n,t}$ specify the model: precisely, $Z^{*}_{t}$ represents the birth probability, indicating the likelihood of encountering a new color (that is a novelty) at time-step $t+1$, while $Q_{n,t}$ signifies the weight assigned to the $n$-th extraction at time-step $t$, reflecting a reinforcement mechanism where the probability of drawing a particular color increases with its past occurrences.
In this framework, the set of possible colors is unbounded, allowing for the continuous introduction of new colors. The urn with infinitely many colors symbolizes the space of all potential events, whereas the sequence of extracted balls represents the realized history. The Poisson-Dirichlet (PD) model James et al. (2008); Pitman (2006); Pitman and Yor (1997) serves as a canonical example of an innovation process. In this model we have
(0.1) $\begin{split}&Z^{*}_{t}=\frac{\theta+\gamma D_{t}}{\theta+t},\qquad Q_{n,t}=\frac{1-\gamma/K_{C_{n},t}}{\theta+t},\\ &\mbox{and so }\quad P_{c,t}=\frac{K_{c,t}-\gamma}{\theta+t},\end{split}$
where $0\leq\gamma<1$, $\theta>-\gamma$, $D_{t}$ denotes the number of distinct colors observed up to time-step $t$ and $K_{c,t}$ is the count of color $c$ up to time-step $t$. This model embodies the “preferential attachment” or “popularity” principle, wherein the probability of observing a particular color is simply proportional to its frequency. The PD process effectively captures empirical laws such as Heaps’ law (Heaps, 1978), which describes the sublinear growth of distinct elements with respect to the total number of observations, and Zipf’s law (Zipf, 1949), characterizing the frequency-rank distribution of elements. Specifically, in the PD model the number of distinct elements $D_{t}$ scales as $D_{t}\propto t^{\gamma}$ for $0<\gamma<1$, and $D_{t}\propto\ln(t)$ when $\gamma=0$.
An intuitive representation of the PD process is the urn with triggering model Aletti and Crimaldi (2021); Tria et al. (2014). An urn initially contains $N_{0}>0$ distinct balls of different colors. Then, at each time-step $t+1$, a ball is drawn at random with replacement and the urn is updated as follows:
  • if the drawn color is new, i.e. not previously observed, then we add $(\widehat{\rho}-1)$ balls of the same color plus $(\nu+1)$ balls of entirely new (not already present in the urn), distinct colors;
  • if the drawn color is old, i.e. it has been already observed, we add $\rho$ balls of the same color.
When the balance condition $\widehat{\rho}+\nu=\rho$ is satisfied, ensuring a constant number of balls added at each time-step, the model reproduces the probabilities specified in equation (0.1) by taking $\rho>\nu\geq 0$ and setting $\theta=N_{0}/\rho$ and $\gamma=\nu/\rho$.
This formulation of the PD process highlights not only the reinforcement mechanism, but also the Kauffman’s principle of the adjacent possible (Kauffman, 2000), wherein the occurrence of a novelty expands the space of potential future novelties. Indeed, the model starts with an urn with a finite number of balls with distinct colors and, whenever a color is extracted for the first time, a set of balls with new colors is added to the urn. This represents Kauffman’s idea that, when a novelty occurs, it triggers further potential novelties.
To explore how different innovation processes influence each other, in Aletti et al. (2023) we have introduced and analyzed a model for a finite network of urns with triggering (see Section 1). Each node in the network possesses its own urn, and the probability of observing a new or old item at a certain node depends on both the node’s own history and the histories of the other nodes. The interactions are governed by two matrices: $\Gamma$ and $W$. Matrix $\Gamma$ modulates the influence of novelties produced in the system on the probability of each node to generate new novelties, extending the Kauffman’s principle of the adjacent possible to the system-level. Matrix $W$ regulates the reinforcement mechanism across nodes, letting the probability of drawing already observed items to be dependent on their frequencies in all the nodes.
When the interaction matrices $\Gamma$ and $W$ are irreducible, the model exhibits synchronization phenomena. Specifically, all nodes produce innovations at the same asymptotic rate, and the distribution of the observed items becomes uniform across the network. These results are formalized in the following Theorems 1.1 and 1.2.
This work aims to:
  • derive the second-order asymptotic behaviour of the model, providing Central Limit Theorems (CLTs) for the quantities of interest;
  • develop statistical tools to make inference on the interaction matrices $\Gamma$ and $W$, as well as on the limiting probabilities of observing an extracted item $c$;
  • apply the theoretical findings to real-world datasets, assessing the interaction intensities in empirical contexts.
It is worthwhile to note that we prove general theoretical results, that can be applied also in other frameworks. Indeed, they regard a pair of multidimensional stochastic processes following a Stochastic Approximation (SA) dynamics (e.g. Laruelle and Pagès (2013); Mokkadem and Pelletier (2006)). However, CLTs from SA literature do not apply to our situation. Indeed, the asymptotic covariance matrices in those results are deterministic, while in our CLTs they are random. This is why we do not use the simple convergence in distribution, but we employ the notion of stable convergence, which is, among other things, essential for obtaining the statistical tools.
0.1. Related literature
While extensive literature exists on interacting urn models with a finite set of colors (see, for instance, Aletti et al. (2017, 2024) and the references therein), studies focusing on urns with infinitely many colors are relatively scarce. Notable exceptions include (Fortini et al., 2018; Iacopini et al., 2020), which explore Dirichlet processes with random reinforcement and urns with triggering, respectively. However, these models differ in key aspects, such as the definition of novelty and the nature of interactions, distinguishing our approach in its system-level application of the adjacent possible principle and the dual reinforcement mechanism. In the model provided in Fortini et al. (2018) (see Example 3.8 in that paper), there is a finite collection of Dirichlet processes with random reinforcement. More precisely, in that model we have a random weight $W_{t,h}$ associated to the extraction at time-step $t$ from the urn $h$ so that, the probability of extracting from urn $h$ an old color (here, the term “old” is not at the system-level, but refers to urn $h$ only) is proportional to the weight associated to that color. The interaction across the urns is introduced by means of the weights, which could be stochastically dependent: each $W_{t,h}$ may be the same for each urn $h$, or a function of the observed outcomes of the other urns, or a function of some common (observable or latent) variables. The present work and Iacopini et al. (2020) share the fact that the considered models are both a collection of urns with triggering with an interacting dynamics that brings the Kauffman’s principle of the adjacent possible from the single agent to the network of agents. Adopting the terminology of Iacopini et al. (2020), we can say that both interacting mechanisms are based on the construction and the updating of a “social” urn for each network node from which the extractions take place, but the construction and the updating rules of the social urns are deeply different in the two models. In particular, differently from Iacopini et al. (2020), we introduce the notion of “new” and “old” at the system level. In Iacopini et al. (2020) the authors focus on the novelties in each sequence (novelty in $h$ = first appearance in $h$ of a new item), that they call “discoveries”; while we also study the sequence of the novelties for the whole system produced by each agent. Furthermore, in Iacopini et al. (2020) the extraction of an “old” item in a certain network node does not affect the other nodes, in our model we also have an interacting reinforcement mechanism for the “old” items: the probability of the extraction of an old item depends on the number of times it has been observed in all the nodes. This allows us to get a specific result on the distribution of the observations in the system among the different items observed.
0.2. Structure of the work
The remainder of the paper is organized as follows. In Section 1 we describe the model and we recall the first-order convergence results. In Section 2 we present the main results of this paper, which consist in the CLTs describing the second-order asymptotic behaviour of the considered system of interacting innovation processes. Then, in Section 3 we illustrate statistical tools derived from the CLTs. To facilitate its application across a wide range of contexts, the inferential methodology is presented within a general framework in the separate appendix, which also includes the general theoretical results and their proofs. Sections, theorems and remarks reported in the appendix are referenced with capital letters. Finally, in Section 4 we apply the proposed methods to two real data sets, evaluating the interaction intensities.
1. Model and first-order convergence results
Suppose to have $N$ urns, labeled from $1$ to $N$. At time-step 0, the colors inside each urn are different from those in the other urns. Let $N_{0,h}>0$ be the number of distinct balls with distinct colors inside the urn $h$. Then, at each time-step $t\geq 1$, one ball is randomly drawn with replacement from each urn and, for any $h=1,\dots N$, urn $h$ is so updated according to the colors extracted from urn $h$ itself and from all the other urns $j\neq h$:
  • if the color $C_{t,h}$ of the ball extracted from urn $h$ is “new” (i.e., it appears for the first time in the system), then into urn $h$ we add $(\widehat{\rho}_{h,h}-1)$, with $\widehat{\rho}_{h,h}>0$, balls of the same color as the extracted one plus $(\nu_{h,h}+1)$, with $\nu_{h,h}\geq 0$, balls of distinct “new” colors (i.e. not already present in the system);
  • if the color $C_{t,h}$ of the ball extracted from urn $h$ is “old” (i.e., it has been already extracted in the system), we add $\rho_{h,h}>0$ balls of the same color into urn $h$;
  • for each $j\neq h$, if the color $C_{t,j}$ of the ball extracted from urn $j$ is “new” (i.e., it appears for the first time in the system), then into urn $h$ we add $\widehat{\rho}_{j,h}\geq 0$ balls of the same color as the one extracted from urn $j$ plus $\nu_{j,h}\geq 0$ balls of distinct “new” colors (i.e. not already present in the system);
  • for each $j\neq h$, if the color $C_{t,j}$ of the ball extracted from urn $j$ is “old” (i.e., it has been already extracted in the system), then into urn $h$ we add $\rho_{j,h}\geq 0$ balls of the same color as the one extracted from urn $j$.
The terms “new” and “old” refer to the entire system, that is a “new” color is a color that has never been extracted from an urn of the system. On the contrary, an “old” color is a color that has already been extracted from at least one urn of the system. We assume that the “new” colors added to a certain urn are always different from those added to the other urns (at the same time-step or in the past). By means of this fact, together with the assumption that initially the colors in the urns are different from each other, we cannot have the same new color extracted simultaneously from different urns. In other words, we cannot have the same novelty observed simultaneously in different processes of the system. Therefore, for each observed new color (novelty) $c$, there exists a unique urn (process), say $j^{*}(c)$, in the system that produced it. However, in a time-step following its first extraction, color $c$ could be also extracted from another urn $h\neq j^{*}(c)$, as a consequence of the interaction among the urns (processes). Indeed, the “contamination” of the color-set of the urn $h$ with the colors present in the other urns is possible by means of the interaction terms $\widehat{\rho}_{j,h}$ and/or $\rho_{j,h}$ in the above model dynamics. Furthermore, as in the standard Poisson-Dirichlet model, we assume the balance condition
[…]
so that, at each time-step, each urn $j$ contributes to increase the number of balls inside urn $h$ by $\rho_{j,h}\geq 0$, with $\rho_{h,h}>0$. Moreover, letting $\rho_{h}=\sum_{j=1}^{N}\rho_{j,h}$ and setting (without loss of generality)
(1.2) $\theta_{h}=N_{0,h}/\rho_{h},\quad\gamma_{j,h}=\nu_{j,h}/\rho_{h},\quad\lambda_{j,h}=\widehat{\rho}_{j,h}/\rho_{h}\quad\mbox{and}\quad w_{j,h}=\rho_{j,h}/\rho_{h}\,$,
we obtain
(1.3) $Z^{*}_{t,h}=P(C_{t+1,h}=\mbox{``new''}\,|\,{\mathbf{C}}_{1},\dots,{\mathbf{C}}_{t})=\frac{\theta_{h}+\sum_{j=1}^{N}\gamma_{j,h}D^{*}_{t,j}}{\theta_{h}+t}\,$,
where ${\mathbf{C}}_{n}=(C_{n,1},\dots,C_{n,N})^{\top}$ with $n=1,\dots,t$, $D^{*}_{t,j}$ denotes the number, until time-step $t$, of distinct observed colors extracted for their first time from urn $j$, that is the number of distinct novelties for the whole system “produced” by urn (process) $j$ until time-step $t$, and, for each “old” color $c$,
[…]
where $K_{t}(j,c)$ denotes the number of times the color $c$ has been extracted from urn $j$ until time-step $t$ and $j^{*}(c)$ denotes the urn from which the color $c$ has been extracted for the first time. Note that the probability that urn (process) $h$ will produce at time-step $t+1$ a novelty for the entire system has an increasing dependence on the number $D^{*}_{t,j}$ of novelties produced by the urn (process) $j$ until time-step $t$ and the parameter $\gamma_{j,h}$ regulates this dependence. In other words, Kauffman’s principle of the adjacent possible is at the “system level”: that is, for each pair $(j,h)$ of urns in the system, the parameter $\gamma_{j,h}$ quantifies how much the production of a novelty by urn $j$ induces potential novelties in urn $h$. Moreover, on the other hand, the probability that from urn $h$ we will extract at time-step $t+1$ an old color $c$ has an increasing dependence on the number $K_{t}(j,c)$ of times the color $c$ has been drawn from urn $j$ until time-step $t$ and the parameter $w_{j,h}$ quantifies how much the number $K_{t}(j,c)$ leads toward a future extraction of a ball of color $c$ from urn $h$. In Aletti et al. (2023) we proved the first-order asymptotic properties of the model:
Theorem 1.1.
Suppose that the matrix $\Gamma=(\gamma_{j,h})_{j,h=1,\dots,N}$ is irreducible and denote by $\gamma^{*}\in(0,1)$ the Perron-Frobenius eigenvalue of $\Gamma$. Then, for each $h=1,\dots,N$, we have
(1.5) $t^{-\gamma^{*}}D^{*}_{t,h}\stackrel{{\scriptstyle a.s.}}{{\longrightarrow}}D^{**}_{\infty,h}\,$,
where $D^{**}_{\infty,h}$ is a finite strictly positive random variable. Moreover, we have
$\frac{D^{*}_{t,h}}{D^{*}_{t,j}}\stackrel{{\scriptstyle a.s.}}{{\longrightarrow}}\frac{u_{h}}{u_{j}}\quad\mbox{and}\quad\frac{D^{*}_{t,h}}{\sum_{j=1}^{N}D^{*}_{t,j}}\stackrel{{\scriptstyle a.s.}}{{\longrightarrow}}\frac{u_{h}}{\sum_{j=1}^{N}u_{j}}\,$,
where ${\mathbf{u}}=(u_{h})_{h}$ is the (up to a multiplicative constant different from zero) left eigenvector of $\Gamma$ associated to $\gamma^{*}$.
Theorem 1.2.
Suppose that the matrix $W=(w_{j,h})_{j,h=1,\dots,N}$ is irreducible. Then, for each $h=1,\dots,N$ and each observed color $c$ in the system, we have
(1.6) $P_{t}(h,c)\stackrel{{\scriptstyle a.s.}}{{\longrightarrow}}\widetilde{P}_{\infty}(c)\,$,
(1.7) $\frac{1}{t}K_{t}(h,c)\stackrel{{\scriptstyle a.s.}}{{\longrightarrow}}\widetilde{P}_{\infty}(c)\quad\mbox{and}\quad\frac{K_{t}(h,c)}{\sum_{j=1}^{N}K_{t}(j,c)}\longrightarrow\frac{1}{N}\,$,
where $\widetilde{P}_{\infty}(c)$ is a random variable that takes values in $(0,1)$ and does not depend on $h$  [^1] Note that in Th 3.2 of Aletti et al. (2023) we state that this random variable takes values in $(0,1]$, but, applying (Aletti et al., 2023, Theorem S1.3) also replacing $P_{t}(h,c)$ by $1-P_{t}(h,c)$, we immediately obtain $(1-\widetilde{P}_{\infty}(c))>0$ a.s., that is $\widetilde{P}_{\infty}(c)<1$ a.s...
The first result states that, if $\Gamma$ is irreducible, that is the graph with the innovation processes as nodes and with $\Gamma$ as the adjacency matrix is strongly connected, then $D^{*}_{t,h}\propto t^{\gamma^{*}}$ a.s. for all $h=1,\dots,N$, that is all the $D^{*}_{t,h}$ grow with the same Heaps’ exponent $\gamma^{*}\in(0,1)$. This means that, at the steady state, all the processes of the network produce innovations for the system at the same rate. In addition, the ratio $D^{*}_{t,h}/D^{*}_{t,j}$ provides a strongly consistent estimator of the ratio $u_{h}/u_{j}$ of the relative eigenvector centrality scores (with respect to $\Gamma^{\top}$) of the two nodes $h$ and $j$ and the share of innovations produced by process $h$ for all the system converges almost surely to the absolute eigenvector centrality score of $h$. The second result affirms that if $W$ is irreducible, that is the graph with the innovation processes as nodes and $W$ as the adjacency matrix is strongly connected, then, for each observed item $c$, the number of times item $c$ has been observed in process $h$ grows linearly. Moreover, the number of times item $c$ has been observed is asymptotically uniformly distributed among the processes of the system and, at the steady state, the probability $\widetilde{P}_{\infty}(c)$ that an item $c$ is observed in a process is the same for each process.
In the next section we complete the analysis of the asymptotics of the model providing some second-order convergence results.
2. Central limit theorems
Assume $N\geq 2$. Let $\Gamma$, $W$, $\Lambda$ be equal to the non-negative $N\times N$ square matrices with elements equal to the model parameters $\gamma_{j,h}$, $w_{j,h}$ (with $w_{h,h}>0$) and $\lambda_{j,h}$ (with $\lambda_{h,h}>0$), respectively. We recall that, by the balance condition (1.1) and the reparametrization (1.2), we have
(2.1) $W=\Gamma+\Lambda,\qquad{{\boldsymbol{1}}}^{\top}\Gamma<{\boldsymbol{1}}^{\top}\qquad\mbox{and}\qquad{{\boldsymbol{1}}}^{\top}W={{\boldsymbol{1}}}^{\top},$
where $\boldsymbol{1}$}} denotes the vector with all the components equal to $1$. As observed above, the matrix $\Gamma$ rules the production of potential novelties and, in particular, its elements out of the diagonal regulate the interaction among the innovation processes with respect to this issue; while, the matrix $W$ rules the interaction among the innovation processes with respect to the choice of an old item. Moreover, set ${\mathbf{D}}^{*}_{t}=(D^{*}_{t,1},\dots,D^{*}_{t,N})^{\top}$, ${\mathbf{D}}^{**}_{\infty}=(D^{**}_{\infty,1},\dots,D^{**}_{\infty,N})^{\top}$, ${\mathbf{P}}_{t}(c)=(P_{t}(1,c),\dots,P_{t}(N,c))^{\top}$ and finally ${\mathbf{K}}_{t}(c)=(K_{t}(1,c),\dots,K_{t}(N,c))^{\top}$.
In order to simplifying the statements of the following theorems, we summarize the required notation and conditions in the following two sets of assumptions:
Assumption 2.1.
Assume that the non-negative matrix $\Gamma=(\gamma_{j,h})_{j,h=1,\dots,N}$ is irreducible. Denote by $\gamma^{*}\in(0,1)$ the Perron-Frobenius eigenvalue of $\Gamma$, by $\mathbf{v}$}} the corresponding right eigenvector with strictly positive entries and such that ${{\mathbf{v}}}^{\top}{{\boldsymbol{1}}}=1$ and denote by $\mathbf{u}$}} the corresponding left eigenvector with strictly positive entries and such that ${{\mathbf{v}}}^{\top}{{\mathbf{u}}}=1$. Finally, let $\gamma^{*}_{2}$ be an eigenvalue of $\Gamma$ different from $\gamma^{*}$ and with highest real part and assume that ${\mathcal{R}e}(\gamma^{*}_{2})/\gamma^{*}<1/2$.
Assumption 2.2.
Assume that the non-negative matrix $W=(w_{j,h})_{j,h=1,\dots,N}$ is irreducible. Denote by ${\mathbf{v}}_{W}$} the right eigenvector associated to the leading eigenvalue $w^{*}=1$, with strictly positive entries and such that ${{\mathbf{v}}_{W}}^{\top}{{\boldsymbol{1}}}=1$. Finally, let $w^{*}_{2}$ be an eigenvalue of $W$ different from $w^{*}=1$ and with highest real part and assume that ${\mathcal{R}e}(w^{*}_{2})<\frac{1}{2}$.
We are now ready to provide the central limit theorems for the considered model, that describe the second-order behaviour of the quantities in Theorems 1.1 and 1.2. The symbol $\boldsymbol{0}$}} denotes the vector with all the components equal to 0.
Theorem 2.1.
Under Assumption 2.1, we have
$t^{\gamma^{*}/2}\left(\frac{{\mathbf{D}}_{t}^{*}}{t^{\gamma^{*}}}-{\mathbf{D}}^{**}_{\infty}\right)\stackrel{{\scriptstyle stably}}{{\longrightarrow}}\mathcal{N}({\boldsymbol{0}},\widetilde{Z}^{**}_{\infty}C_{det,\Gamma})\,$,
where $\widetilde{Z}_{\infty}^{**}$ is an integrable strictly positive random variable such that ${\mathbf{D}}^{**}_{\infty}=\frac{\widetilde{Z}^{**}_{\infty}}{\gamma^{*}}\,{\mathbf{u}}$ and $C_{det,\Gamma}$ is a deterministic matrix depending on the eigen-structure of $\Gamma$, which is possible to write explicitly.
Proof.
Denote by $X^{*}_{t,h}$ the random variable that takes value $1$ when the ball extracted from urn $h$ at time-step $t$ has a new (for all the system) color and is equal to 0 otherwise. Then $Z^{*}_{t,h}$ coincides with $P(X^{*}_{t+1,h}=1\,|\,{\mathbf{C}}_{1},\dots,{\mathbf{C}}_{t})=E[X^{*}_{t+1,h}\,|\,{\mathbf{C}}_{1},\dots,{\mathbf{C}}_{t}]$ and $D^{*}_{t,h}$ can be written as $\sum_{n=1}^{t}X^{*}_{n,h}$. Since we have
$Z^{*}_{t,h}=\frac{\theta_{h}+\sum_{n=1}^{t}\sum_{j=1}^{N}\gamma_{j,h}X^{*}_{n,j}}{\theta_{h}+t}\,$,
we obtain the following dynamics for $Z^{*}_{t,h}$:
$Z^{*}_{0,h}=1,\qquad Z^{*}_{t+1,h}=(1-r_{t,h})Z^{*}_{t,h}+r_{t,h}\sum_{j=1}^{N}\gamma_{j,h}X^{*}_{t+1,j}\quad\mbox{for }t\geq 0\,$,
where $r_{t,h}=1/(\theta_{h}+t+1)=1/(t+1)+O(1/t^{2})$  [^2] In all the paper, the notation $O(s_{t})$ denotes a generic (possibly random) remainder term $R_{t}$ such that $|R_{t}|\leq Cs_{t}$ for a suitable constant $C$ and for $t$ large enough.. The corresponding vectorial dynamics for ${\mathbf{Z}}^{*}_{t}=(Z^{*}_{t,1},\dots,Z^{*}_{t,N})^{\top}$ is
(2.2) $\begin{split}{{\mathbf{Z}}}^{*}_{0}&={\boldsymbol{1}}\\ {{\mathbf{Z}}}^{*}_{t+1}&=\left(1-\frac{1}{t+1}\right){{\mathbf{Z}}}^{*}_{t}+\frac{1}{t+1}\Gamma^{\top}{{\mathbf{X}}}_{t+1}^{*}+O(1/t^{2}){\boldsymbol{1}}\\ &=\left[1-\frac{1}{t+1}(I-\Gamma^{\top})\right]{{\mathbf{Z}}}^{*}_{t}+\frac{1}{t+1}\Gamma^{\top}\Delta{{\mathbf{M}}}^{*}_{t+1}+O(1/t^{2}){\boldsymbol{1}}\quad\mbox{for }t\geq 0,\end{split}$
where $\Delta{{\mathbf{M}}}^{*}_{t+1}={{\mathbf{X}}}^{*}_{t+1}-{{\mathbf{Z}}}^{*}_{t}$. Now, we fix $x>0$ and set
$\zeta_{0}(x)=1,\qquad\zeta_{t}(x)=\frac{\Gamma(t+x)}{\Gamma(t)}\sim t^{x}\uparrow+\infty\,.$
More precisely, from (Gouet, 1993, Lemma 4.1) we have
$\zeta_{t}(x)=t^{x}+O(t^{x-1})\quad\mbox{and}\quad\frac{1}{\zeta_{t}(x)}=\frac{1}{t^{x}}+O(1/t^{x+1}),$
and so
(2.3) $\begin{split}\frac{1}{\zeta_{t+1}(x)\zeta_{t+1}(1-x)}&=\left(\frac{1}{(t+1)^{x}}+O(1/t^{x+1})\right)\left(\frac{1}{(t+1)^{1-x}}+O(1/t^{2-x})\right)\\ &=\frac{1}{t+1}+O(1/t^{2}).\end{split}$
Hence, multiplying (2.2) by $\zeta_{t+1}(1-\gamma^{*})$ and using the relation
$\frac{\zeta_{t+1}(x)}{\zeta_{t}(x)}=\frac{\Gamma(t+x+1)}{\Gamma(t+1)}\frac{\Gamma(t)}{\Gamma(t+x)}=1+\frac{x}{t}=1+\frac{x}{t+1}+O(1/t^{2}),$
with $x=1-\gamma^{*}$, we get the following dynamics for ${{\mathbf{Z}}}^{**}_{t}=\zeta_{t}(1-\gamma^{*}){{\mathbf{Z}}}^{*}_{t}$, where we set $\Delta{{\mathbf{M}}}^{**}_{t+1}=\zeta_{t}(1-\gamma^{*})\Delta{{\mathbf{M}}}^{*}_{t+1}$:
$\begin{split}{{\mathbf{Z}}}^{**}_{0}&={\boldsymbol{1}}\\ {{\mathbf{Z}}}^{**}_{t+1}&=\left[1-\frac{1}{t+1}(I-\Gamma^{\top})\right]\frac{\zeta_{t+1}(1-\gamma^{*})}{\zeta_{t}(1-\gamma^{*})}{{\mathbf{Z}}}^{**}_{t}+\frac{1}{t+1}\Gamma^{\top}\Delta{{\mathbf{M}}}^{**}_{t+1}+O\left(\frac{\zeta_{t+1}(1-\gamma^{*})}{t^{2}}\right){\boldsymbol{1}}\\ &=\left[1-\frac{1}{t+1}(I-\Gamma^{\top})\right]\left(1+\frac{1-\gamma^{*}}{t+1}+O(1/t^{2})\right){{\mathbf{Z}}}^{**}_{t}+\frac{1}{t+1}\Gamma^{\top}\Delta{{\mathbf{M}}}^{**}_{t+1}+O\left(\frac{\zeta_{t+1}(1-\gamma^{*})}{t^{2}}\right){\boldsymbol{1}}\\ &={{\mathbf{Z}}}^{**}_{t}-\frac{1}{t+1}(\gamma^{*}I-\Gamma^{\top}){{\mathbf{Z}}}^{**}_{t}+\frac{1}{t+1}\Gamma^{\top}\Delta{{\mathbf{M}}}^{**}_{t+1}+O\left(\frac{\zeta_{t+1}(1-\gamma^{*})}{t^{2}}\right){\boldsymbol{1}}\,.\end{split}$
Moreover, recalling that $D^{*}_{t,h}=\sum_{n=1}^{t}X^{*}_{n,h}$, we find
$\begin{split}\frac{{\mathbf{D}}^{*}_{0}}{\zeta_{0}(\gamma^{*})}&={\boldsymbol{0}}\\ \frac{{\mathbf{D}}^{*}_{t+1}}{\zeta_{t+1}(\gamma^{*})}&=\frac{\zeta_{t}(\gamma^{*})}{\zeta_{t+1}(\gamma^{*})}\frac{{\mathbf{D}}_{t}^{*}}{\zeta_{t}(\gamma^{*})}+\frac{1}{\zeta_{t+1}(\gamma^{*})}{\mathbf{X}}^{*}_{t+1}\\ &=\left(1-\frac{\gamma^{*}}{t+1}\right)\frac{{\mathbf{D}}_{t}^{*}}{\zeta_{t}(\gamma^{*})}+\frac{1}{\zeta_{t+1}(\gamma^{*})}\Delta{\mathbf{M}}^{*}_{t+1}+\frac{1}{\zeta_{t+1}(\gamma^{*})}{\mathbf{Z}}^{*}_{t}\\ &=\left(1-\frac{\gamma^{*}}{t+1}\right)\frac{{\mathbf{D}}_{t}^{*}}{\zeta_{t}(\gamma^{*})}+\frac{1}{\zeta_{t+1}(\gamma^{*})\zeta_{t+1}(1-\gamma^{*})}\Delta{\mathbf{M}}^{**}_{t+1}\\ &+\frac{1}{\zeta_{t+1}(\gamma^{*})\zeta_{t+1}(1-\gamma^{*})}\frac{\zeta_{t+1}(1-\gamma^{*})}{\zeta_{t}(1-\gamma^{*})}{\mathbf{Z}}^{**}_{t}\,.\end{split}$
Using (2.3) and the relation $\zeta_{t+1}(x)/\zeta_{t}(x)=1+O(1/t)$, and combining together with the previous dynamics of $({{\mathbf{Z}}}^{**}_{t})_{t}$ and $({\mathbf{D}}^{*}_{t+1}/\zeta_{t+1}(\gamma^{*}))_{t}$, we obtain
[…]
where $O(\zeta_{t}(1-\gamma^{*})/t^{2})=O(1/t^{1+\gamma^{*}})$. Moreover, we recall that in (Aletti et al., 2023, Theorem S1.1 and Lemma S1.2) we proved
[…]
where $\widetilde{Z}_{\infty}^{**}$ is an integrable strictly positive random variable and ${\mathbf{D}}^{**}_{\infty}=\frac{\widetilde{Z}^{**}_{\infty}}{\gamma^{*}}\,{\mathbf{u}}$. Therefore, we can apply Theorem A.4 with ${\mathbf{A}}_{t}={\mathbf{Z}}^{**}_{t}=\zeta_{t}(1-\gamma^{*}){\mathbf{Z}}^{*}_{t}$, ${\mathbf{B}}_{t}=\frac{{\mathbf{D}}_{t}^{*}}{\zeta_{t}(\gamma^{*})}$, $\Phi=\Gamma$, $\phi^{*}=\gamma^{*}$, $\Sigma_{\infty}=\widetilde{Z}^{**}_{\infty}diag({\mathbf{u}})$, $\Delta{{\mathbf{M}}}_{t+1}=\Delta{\mathbf{M}}^{**}_{t+1}=\zeta_{t}(1-\gamma^{*})\Delta{\mathbf{M}}^{*}_{t+1}=\zeta_{t}(1-\gamma^{*})({\mathbf{X}}^{*}_{t+1}-{\mathbf{Z}}^{*}_{t})$. Finally, we can conclude by observing that, since $\gamma^{*}<1$, we have
$t^{\gamma^{*}/2}\left(\frac{{\mathbf{D}}_{t}^{*}}{t^{\gamma^{*}}}-\frac{{\mathbf{D}}_{t}^{*}}{\zeta_{t}(\gamma^{*})}\right)=t^{\gamma^{*}/2}{\mathbf{D}}_{t}^{*}O\left(1/t^{\gamma^{*}+1}\right)=\frac{{\mathbf{D}}^{*}_{t}}{t^{\gamma^{*}}}O\left(1/t^{1-\gamma^{*}/2}\right)\stackrel{{\scriptstyle a.s.}}{{\longrightarrow}}{\boldsymbol{0}}\,.$
Note that in Theorem A.4 we have the explicit formulas for the elements of the matrix $C_{det,\Gamma}$: that is, $C_{det,\Gamma}=1/\gamma^{*}$ when $N=1$ and
$C_{det,\Gamma}=\frac{1}{\gamma^{*}}({\mathbf{v}}^{\top}diag({\mathbf{u}}){\mathbf{v}}){\mathbf{u}}{\mathbf{u}}^{\top}+\frac{1}{\widetilde{Z}^{**}_{\infty}}{\mathcal{M}}^{33}_{\infty}\,$,
where ${\mathcal{M}}^{33}_{\infty}$ is the matrix defined in Theorem A.3 when $N\geq 2$ with the matrix $\Phi$ equal to $\Gamma$. Note that, since the expression for ${\mathcal{M}}^{33}_{\infty}$ contains the matrix $\Sigma_{\infty}$, and so the term $\widetilde{Z}^{**}_{\infty}$, we have that the matrix $\frac{1}{\widetilde{Z}^{**}_{\infty}}{\mathcal{M}}^{33}_{\infty}$ is deterministic. ∎
Theorem 2.2.
Under Assumption 2.1, we have
$t^{\gamma^{*}/2}\left(\frac{{\mathbf{D}}_{t}^{*}}{\sum_{h=1}^{N}D^{*}_{t,h}}-\frac{{\mathbf{u}}}{\sum_{h=1}^{N}u_{h}}\right)\stackrel{{\scriptstyle stably}}{{\longrightarrow}}\mathcal{N}\left({\boldsymbol{0}},\frac{1}{\widetilde{Z}^{**}_{\infty}}\frac{(\gamma^{*})^{2}}{(\sum_{j=1}^{N}u_{j})^{2}}Q_{\Gamma}C_{det,\Gamma}Q_{\Gamma}^{\top}\right)\,$,
and
$\sqrt{\sum_{h=1}^{N}D^{*}_{t,h}}\left(\frac{{\mathbf{D}}_{t}^{*}}{\sum_{h=1}^{N}D^{*}_{t,h}}-\frac{{\mathbf{u}}}{\sum_{h=1}^{N}u_{h}}\right)\stackrel{{\scriptstyle stably}}{{\longrightarrow}}\mathcal{N}\left({\boldsymbol{0}},\frac{\gamma^{*}}{\sum_{j=1}^{N}u_{j}}Q_{\Gamma}C_{det,\Gamma}Q_{\Gamma}^{\top}\right)\,$,
where $Q_{\Gamma}=I-{\mathbf{u}}{\boldsymbol{1}}^{\top}/{\boldsymbol{1}}^{\top}{\mathbf{u}}$ and $C_{det,\Gamma}$ is defined in Theorem 2.1. Moreover, for each $j$ we have
$t^{\gamma^{*}/2}\left(\frac{{\mathbf{D}}_{t}^{*}}{D^{*}_{t,j}}-\frac{{\mathbf{u}}}{u_{j}}\right)\stackrel{{\scriptstyle stably}}{{\longrightarrow}}\mathcal{N}\left({\boldsymbol{0}},\frac{1}{\widetilde{Z}^{**}_{\infty}}\frac{(\gamma^{*})^{2}}{u_{j}^{2}}Q_{j}C_{det,\Gamma}Q_{j}^{\top}\right)\,$,
and
$\sqrt{D^{*}_{t,j}}\left(\frac{{\mathbf{D}}_{t}^{*}}{D^{*}_{t,j}}-\frac{{\mathbf{u}}}{u_{j}}\right)\stackrel{{\scriptstyle stably}}{{\longrightarrow}}\mathcal{N}\left({\boldsymbol{0}},\frac{\gamma^{*}}{u_{j}}Q_{j}C_{det,\Gamma}Q_{j}^{\top}\right)\,$,
where $Q_{j}=I-{\mathbf{u}}{\mathbf{e}}_{j}^{\top}/u_{j}$ with ${\mathbf{e}}_{j}$ the $j$th vector of the canonical base and $C_{det,\Gamma}$ is defined in Theorem 2.1.
Proof.
We have to apply Corollary A.6 reported in the appendix with ${\mathbf{A}}_{t}={\mathbf{Z}}^{**}_{t}=\zeta_{t}(1-\gamma^{*}){\mathbf{Z}}^{*}_{t}$, ${\mathbf{B}}_{t}=\frac{{\mathbf{D}}_{t}^{*}}{\zeta_{t}(\gamma^{*})}$, ${\mathbf{x}}={\boldsymbol{1}}$, $\widetilde{A}_{\infty}=\widetilde{Z}_{\infty}^{**}$ and $\phi^{*}=\gamma^{*}$. Note that, as specified in the proof of this corollary, we have $\widetilde{Z}_{\infty}^{**}Q_{\Gamma}C_{det,\Gamma}Q_{\Gamma}^{\top}=Q_{\Gamma}{\mathcal{M}}^{33}_{\infty}Q_{\Gamma}^{\top}$, where ${\mathcal{M}}^{33}_{\infty}$ is the matrix defined in Theorem A.3] when $N\geq 2$ with the matrix $\Phi$ equal to $\Gamma$. Similarly, we obtain the second part of the statement by means of Corollary A.6 with ${\mathbf{x}}={\mathbf{e}}_{j}$. ∎
Theorem 2.3.
Under Assumption 2.2, we have for each observed color $c$ in the system
$\sqrt{t}\left(\frac{{\mathbf{K}}_{t}(c)}{t}-\widetilde{P}_{\infty}(c){\boldsymbol{1}}\right)\stackrel{{\scriptstyle stably}}{{\longrightarrow}}\mathcal{N}({\boldsymbol{0}},\widetilde{P}_{\infty}(c)(1-\widetilde{P}_{\infty}(c))C_{det,W})\,$,
where $\widetilde{P}_{\infty}(c)=a.s.-\lim_{t}P_{t}(h,c)=a.s.-\lim_{t}K_{t}(h,c)/t$ for each $h=1,\dots,N$, $\widetilde{P}_{\infty}(c)\in(0,1)$ a.s. and $C_{det,W}$ is a deterministic matrix depending on the eigen-structure of $W$, which is possible to write explicitly.
Proof.
For any color $c$ already observed in the system at time $t$, we can write
$P_{t}(h,c)=\frac{\sum_{n=1}^{t}\sum_{j=1}^{N}w_{j,h}\Delta K_{n}(j,c)}{\theta_{h}+t}-\frac{\gamma_{j^{*}(c),h}}{{\theta_{h}+t}}\,$,
where $\Delta K_{n}(j,c)=K_{n}(j,c)-K_{n-1}(j,c)$ takes values in $\{0,1\}$ and $E[\Delta K_{n+1}(j,c)|\mbox{past}]=P_{n}(j,c)$. Then, we obtain the following dynamics for $P_{t}(h,c)$:
$P_{t+1}(h,c)=(1-r_{t,h})P_{t}(h,c)+r_{t,h}\sum_{j=1}^{N}w_{j,h}\Delta K_{t+1}(j,c)\,$,
where $r_{t,h}=1/(\theta_{h}+t+1)=1/(t+1)+O(1/t^{2})$. Thus the corresponding vectorial dynamics for ${\mathbf{P}}_{t}(c)=(P_{t}(1,c),\dots,P_{t}(N,c))^{\top}$ is
$\begin{split}{\mathbf{P}}_{t^{*}(c)}(c)&\neq{\boldsymbol{0}},\\ {\mathbf{P}}_{t+1}(c)&=\left(1-\frac{1}{t+1}\right){\mathbf{P}}_{t}(c)+\frac{1}{t+1}W^{T}{\Delta{\mathbf{K}}}_{t+1}(c)+O(1/t^{2}){\boldsymbol{1}}\\ &={\mathbf{P}}_{t}(c)-\frac{1}{t+1}(I-W^{\top}){\mathbf{P}}_{t}(c)+\frac{1}{t+1}W^{\top}\Delta{{\mathbf{M}}}_{t+1}(c)+O(1/t^{2}){\boldsymbol{1}},\end{split}$
for $t\geq t^{*}(c)$, where $t^{*}(c)$ is the time-step of the first extraction of $c$ in the system, ${\Delta{\mathbf{K}}}_{t}(c)=(\Delta K_{t}(1,c),\dots,\Delta K_{t}(N,c))^{\top}$ and $\Delta{{\mathbf{M}}}_{t+1}(c)={\Delta{\mathbf{K}}}_{t+1}(c)-{\mathbf{P}}_{t}(c)$.
Considering also the dynamics of ${\mathbf{K}}_{t}(c)/t$, we obtain for $t\geq t^{*}(c)$
[…]
Moreover, from Theorem 1.2, we have $\widetilde{P}_{t}(c)={\mathbf{v}}_{W}^{\top}{\mathbf{P}}_{t}(c)\stackrel{{\scriptstyle a.s.}}{{\longrightarrow}}\widetilde{P}_{\infty}(c)\,$, and this convergence is also in mean because $(\widetilde{P}_{t}(c))_{t}$ is uniformly bounded. Therefore, we can apply Theorem A.4 with ${\mathbf{A}}_{t}={\mathbf{P}}_{t}(c)$, ${\mathbf{B}}_{t}={\mathbf{K}}_{t}(c)/t$, $\Phi=W$, $\phi^{*}=w^{*}=1$, ${\mathbf{u}}={\mathbf{u}}_{W}={\boldsymbol{1}}$, ${\mathbf{v}}={\mathbf{v}}_{W}$, $\Sigma_{\infty}=\widetilde{P}_{\infty}(c)(1-\widetilde{P}_{\infty}(c))I$, $\Delta{{\mathbf{M}}}_{t+1}=\Delta{{\mathbf{M}}}_{t+1}(c)={\Delta{\mathbf{K}}}_{t+1}(c)-{\mathbf{P}}_{t}(c)$. Note that in Theorem A.4 we have the explicit formulas for the elements of the matrix $C_{det,W}$: that is, $C_{det,W}=1$ when $N=1$ and
$C_{det,W}=({\mathbf{v}}_{W}^{\top}{\mathbf{v}}_{W}){\boldsymbol{1}}{\boldsymbol{1}}^{\top}+\frac{1}{\widetilde{P}_{\infty}(c)(1-\widetilde{P}_{\infty}(c))}{\mathcal{M}}^{33}_{\infty}\,$,
where ${\mathcal{M}}^{33}_{\infty}$ is the matrix defined in Theorem A.3] when $N\geq 2$ with the matrix $\Phi$ equal to $W$. Note that, since the expression for ${\mathcal{M}}^{33}_{\infty}$ contains the matrix $\Sigma_{\infty}$, and so the factor $\widetilde{P}_{\infty}(c)(1-\widetilde{P}_{\infty}(c))$, we have that the matrix $\frac{1}{\widetilde{P}_{\infty}(c)(1-\widetilde{P}_{\infty}(c))}{\mathcal{M}}^{33}_{\infty}$ is deterministic. ∎
Theorem 2.4.
Under Assumption 2.2, we have for each observed color $c$ in the system
$\sqrt{t}\left(\frac{{\mathbf{K}}_{t}(c)}{\sum_{h=1}^{N}K_{t}(h,c)}-\frac{1}{N}{\boldsymbol{1}}\right)\stackrel{{\scriptstyle stably}}{{\longrightarrow}}{\mathcal{N}}\left({\boldsymbol{0}},\frac{(1-\widetilde{P}_{\infty}(c))}{N^{2}\widetilde{P}_{\infty}(c)}Q_{W}C_{det,W}Q_{W}^{\top}\right),$
and
$\sqrt{\frac{\sum_{h=1}^{N}K_{t}(h,c)}{1-(Nt)^{-1}\sum_{h=1}^{N}K_{t}(h,c)}}\left(\frac{{\mathbf{K}}_{t}(c)}{\sum_{h=1}^{N}K_{t}(h,c)}-\frac{1}{N}{\boldsymbol{1}}\right)\stackrel{{\scriptstyle stably}}{{\longrightarrow}}{\mathcal{N}}\left({\boldsymbol{0}},\frac{1}{N}Q_{W}C_{det,W}Q_{W}^{\top}\right).$
where $Q_{W}=(I-N^{-1}{\boldsymbol{1}}{\boldsymbol{1}}^{\top})$ and $C_{det,W}$ is defined in Theorem 2.3.
Proof.
We have to apply Corollary A.6 reported in the appendix with ${\mathbf{A}}_{t}={\mathbf{P}}_{t}(c)$, ${\mathbf{B}}_{t}={\mathbf{K}}_{t}(c)/t$, ${\mathbf{x}}={\boldsymbol{1}}$, $\widetilde{A}_{\infty}=\widetilde{P}_{\infty}(c)$, $\phi^{*}=w^{*}=1$ and ${\mathbf{u}}={\mathbf{u}}_{W}={\boldsymbol{1}}$. Note that, as specified in the proof of this corollary, we have $\widetilde{P}_{\infty}(c)(1-\widetilde{P}_{\infty}(c))Q_{W}C_{det,W}Q_{W}^{\top}=Q_{W}{\mathcal{M}}^{33}_{\infty}Q_{W}^{\top}$, where ${\mathcal{M}}^{33}_{\infty}$ is the matrix defined in Theorem A.3 with the matrix $\Phi$ equal to $W$. ∎
Remark 2.1.
When ${\mathcal{R}e}(\gamma^{*}_{2})/\gamma^{*}\geq 1/2$, as highlighted in Remark A.1, the rate of convergence is slower than $t^{\gamma^{*}/2}$. Indeed, similarly to the stochastic approximation literature (e.g. Zhang (2016)), although we deal with a random limit covariance matrix, for the case ${\mathcal{R}e}(\gamma^{*}_{2})/\gamma^{*}=1/2$ we conjecture a CLT with a logarithmic correction in the scaling; while in the case ${\mathcal{R}e}(\gamma^{*}_{2})/\gamma^{*}>1/2$ we expect a power-law rate with an exponent smaller than $\gamma^{*}/2$, but not a Gaussian limit. Analogous considerations can be done for the cases ${\mathcal{R}e}(w^{*}_{2})=1/2$ and ${\mathcal{R}e}(w^{*}_{2})>1/2$.
3. Statistical inference
We here present some statistical tools based on the results of Section 2. More precisely, we provide a statistical test for each of the two interaction matrices, $\Gamma$ and $W$, based on the observable processes ${\mathbf{D}}^{*}_{t}$ and ${\mathbf{K}}_{t}(c)$. Furthermore, we use the latter process to provide also a confidence interval for the limit probability $\widetilde{P}_{\infty}(c)$. The general procedure to be followed in order to obtain these statistical instruments is described in Sections B C. In particular, in our framework, we have to apply it with ${\mathbf{B}}_{t}=\frac{{\mathbf{D}}_{t}^{*}}{\zeta_{t}(\gamma^{*})}$ (or, equivalently ${\mathbf{B}}_{t}=\frac{{\mathbf{D}}_{t}^{*}}{t^{\gamma^{*}}}$) and with ${\mathbf{B}}_{t}={\mathbf{K}}_{t}(c)/t$. To this end, we observe that Assumption B.1 is true in both cases with $\Sigma_{det}=\Sigma_{det,\Gamma}=diag({\mathbf{u}})$ in the first one and $\Sigma_{det}=\Sigma_{det,W}=I$ in the second one.
We firstly focus on the case $N=2$ and, then, we consider a particular type of interaction (mean-field interaction) in the general case $N\geq 2$. It is worthwhile to underline that the statistical tools described in Section B are very general and so it is also possible to explicit the formulas for any family of interaction matrices.
3.1. The case $N=2$
3.1.1. Hypothesis test on $\Gamma$
In what follows we assume the parameters $\gamma^{*}\in(0,1)$ and $r=u_{1}/u_{2}>0$ to be known.
We use the re-parametrization of the matrix $\Gamma$ given by the constrains
$r\gamma_{1,1}+\gamma_{2,1}=r\gamma^{*}\quad\mbox{and}\quad r\gamma_{1,2}+\gamma_{2,2}=\gamma^{*}\,.$
and make inference on the other two parameters $\gamma_{1,2}$ and $\gamma_{2,1}$ (that have to be such that the assumptions on the matrix $\Gamma$ are verified). If we use the parametrization $\iota=\gamma_{1,2}+\gamma_{2,1}$ and $\eta=\tfrac{\gamma_{1,2}}{\gamma_{1,2}+\gamma_{2,1}}$ so that $\iota$ rules the intensity of the whole interaction between the two processes due to matrix $\Gamma$, while $\eta$ (resp. $(1-\eta)$) denotes the percentage of this intensity due to the influence of process $1$ on process $2$ (resp., of process $2$ on process $1$), the expressions for the elements of the matrix $\Gamma$ become
(3.1) $\gamma_{1,1}=\gamma^{*}-\frac{(1-\eta)\iota}{r},\quad\gamma_{1,2}=\eta\iota,\quad\gamma_{2,1}=(1-\eta)\iota\quad\mbox{and}\quad\gamma_{2,2}=\gamma^{*}-r\eta\iota,$
where (by the required conditions on $\Gamma$)
(3.2) $\eta\in(0,1)\quad\mbox{and}\quad\iota\in J_{\eta}\quad\mbox{with},$
$J_{\eta}=\left(0,\min\big{(}\tfrac{\gamma^{*}r}{1-\eta},\tfrac{\gamma^{*}}{r\eta}\big{)}\right]\cap\left(0,\tfrac{1-\gamma^{*}}{\eta(1-r)}I_{(r<1)}+\tfrac{1-\gamma^{*}}{(1-\eta)(1-1/r)}I_{(r>1)}+\infty I_{(r=1)}\right).$
Two-sided test
If we take the null hypothesis $H_{0}:\;\iota=\iota_{0}\,,\,\eta=\eta_{0}$, where $\iota_{0}$ and $\eta_{0}$ satisfy (3.2), i.e. $\iota_{0}\in J_{\eta_{0}}$, and the required condition in Assumption 2.1 for the second eigenvalue of $\Gamma$, that is $\iota_{0}\frac{(1-\eta_{0})+r^{2}\eta_{0}}{r}>\gamma^{*}/2$, then we can use Theorem 2.1 and obtain the test statistic (see (C.1) from Remark C.1 with $\phi=\gamma^{*}$, $\phi_{2}^{*}=\gamma^{*}_{2}=\gamma^{*}-[(1-\eta_{0})\iota_{0}+r^{2}\eta_{0}\iota_{0}]/r$, $g(x)=x$, and ${\mathbf{B}}_{t}=\frac{{\mathbf{D}}_{t}^{*}}{\zeta_{t}(\gamma^{*})}$ or, equivalently ${\mathbf{B}}_{t}=\frac{{\mathbf{D}}_{t}^{*}}{t^{\gamma^{*}}}$)
(3.3) $q_{r,\eta_{0}}\Delta_{0}\,\frac{(D^{*}_{1,t}-rD^{*}_{2,t})^{2}}{\widetilde{D}^{*}_{t}}\stackrel{{\scriptstyle d}}{{\longrightarrow}}\chi^{2}(1)\quad\mbox{under }H_{0}\,$,
where $\widetilde{D}^{*}_{t}={\mathbf{v}}^{\top}{\mathbf{D}}^{*}_{t}=\frac{\eta_{0}rD^{*}_{1,t}+(1-\eta_{0})D^{*}_{2,t}}{\eta_{0}r+(1-\eta_{0})}$, $q_{r,\eta_{0}}=\frac{2[(1-\eta_{0})+\eta_{0}r^{2}]}{r(1+r)[(1-\eta_{0})+\eta_{0}r]}$ and
$\Delta_{0}=\frac{1}{2}-\frac{\gamma^{*}_{2}}{\gamma^{*}}=\frac{\iota_{0}}{\gamma^{*}}\frac{(1-\eta_{0})+\eta_{0}r^{2}}{r}-\frac{1}{2}\,.$
It is immediate to see that the test statistic (3.3) has an increasing dependence on the parameter $\iota_{0}$, while its behaviour with respect the parameter $\eta_{0}$ is not clear. However, we can observe that it can be rewritten as
$\frac{2}{r(1+r)}\frac{(1-\eta_{0})+\eta_{0}r^{2}}{(1-\eta_{0})+\eta_{0}rD^{*}_{1,t}/D^{*}_{2,t}}\Delta_{0}D^{*}_{2,t}(D^{*}_{1,t}/D^{*}_{2,t}-r)^{2}\,$,
where, for any value of $\eta_{0}$, we have
$\frac{(1-\eta_{0})+\eta_{0}r^{2}}{(1-\eta_{0})+\eta_{0}rD^{*}_{1,t}/D^{*}_{2,t}}\stackrel{{\scriptstyle a.s.}}{{\longrightarrow}}1\,.$
Hence, we can simplify (3.3) as
(3.4) $\tfrac{2}{r(1+r)}\Delta_{0}D^{*}_{2,t}(D^{*}_{1,t}/D^{*}_{2,t}-r)^{2}\stackrel{{\scriptstyle d}}{{\longrightarrow}}\chi^{2}(1)\quad\mbox{under }H_{0}\,$,
where the dependence on $\iota_{0}$ and $\eta_{0}$ is only in the factor $\Delta_{0}$. Notice also that in the present framework the two eigenvalues $\gamma^{*}$ and $\gamma_{2}^{*}$ are both real and so the quantity $\Delta_{0}$ represents the distance of the ratio between the second and the leading eigenvalue from the critical value $1/2$ (see condition ${\mathcal{R}e}(\gamma^{*}_{2})/\gamma^{*}<1/2$ in Assumption 2.1). Moreover, $\Delta_{0}$ is the only quantity in the test statistics (3.4) that contains the parameters $\iota_{0}$ and $\eta_{0}$. Moreover, it is neither random nor time-dependent. The random quantity $D^{*}_{2,t}(D^{*}_{1,t}/D^{*}_{2,t}-r)^{2}=[\sqrt{D^{*}_{2,t}}(D^{*}_{1,t}/D^{*}_{2,t}-r)]^{2}$ describes the fluctuations of $D^{*}_{1,t}/D^{*}_{2,t}$ around $r$ (note that (3.4) is coherent with the convergence result stated in Theorem 2.2).
One-sided test
Since $\Delta_{0}$, and so the statistic (3.4), is increasing in $\iota_{0}$, we can replace the condition $\iota=\iota_{0}$ in $H_{0}$ by $\iota\geq\iota_{0}$. Further, the quantity $\Delta_{0}$ is increasing in $\eta_{0}$ when $r>1$ and decreasing in $\eta_{0}$ when $r<1$. Hence, the statistic (3.4) works well also if we replace the condition $\eta=\eta_{0}$ in $H_{0}$ by $\eta\geq\eta_{0}$ when $r>1$ or by $\eta\leq\eta_{0}$ when $r<1$. Therefore, if we set $\eta_{0}=1/2$, we get
$\Delta_{0}=\frac{1}{2}-\frac{\gamma^{*}_{2}}{\gamma^{*}}=\frac{\iota_{0}}{\gamma^{*}}\frac{1+r^{2}}{2r}-\frac{1}{2}$
and we can perform a one-sided test with one of the following two null hypotheses, according to the value of $r$:
  • Case $r>1$: $H_{0}:\;\iota\geq\iota_{0},\,\eta\geq 1/2$ (the interaction intensity is at least $\iota_{0}$ and the processes equally influence each other or the interaction is mostly due to the influence of process $1$ on process $2$), where $\iota_{0}$ satisfies (3.2) for some $\eta$ in $H_{0}$, i.e. $\iota_{0}\in\cup_{\eta\geq 1/2}J_{\eta}$, and $\iota_{0}>\gamma^{*}\frac{r}{1+r^{2}}$;
  • Case $r<1$: $H_{0}:\;\iota\geq\iota_{0},\,\eta\leq 1/2$ (the interaction intensity is at least $\iota_{0}$ and the processes equally influence each other or the interaction is mostly due to the influence of process $2$ on process $1$), where $\iota_{0}$ satisfies (3.2) for some $\eta$ in $H_{0}$, i.e. $\iota_{0}\in\cup_{\eta\leq 1/2}J_{\eta}$, and $\iota_{0}>\gamma^{*}\frac{r}{1+r^{2}}$.
Finally, in the very special case when $r=1$, the test statistic (3.4) does not depend on the value of $\eta_{0}$ and it can be used for testing the intensity $\iota$ only.
Power
Following Remark B.1, we find that, under the alternative $H_{1}:\,\iota=\iota_{1},\,\eta=\eta_{1}$, with $\iota_{1}\in J_{\eta_{1}}$ and $\iota_{1}\frac{(1-\eta_{1})+r^{2}\eta_{1}}{r}>\gamma^{*}/2$, the statistic (3.4) is asymptotically distributed as
$\frac{\Delta_{0}}{\Delta_{1}}\,\chi^{2}(1)=\frac{\frac{\iota_{0}}{\gamma^{*}}\frac{(1-\eta_{0})+\eta_{0}r^{2}}{r}-\frac{1}{2}}{\frac{\iota_{1}}{\gamma^{*}}\frac{(1-\eta_{1})+\eta_{1}r^{2}}{r}-\frac{1}{2}}\,\chi^{2}(1)\,.$
Notice that the term $\tfrac{\Delta_{0}}{\Delta_{1}}$ is time-invariant and this means that increasing the time $t$ does not increase the power. Nevertheless, the power increases as $\eta_{1}$ decreases if $r>1$ or as $\eta_{1}$ increases if $r<1$. In addition, the power improves as the value of $\iota_{1}$ decreases and it tends toward $1$ as $\Delta_{1}\to 0$, that is as $\iota_{1}$ approaches its minimum admissible value $\tfrac{r}{(1-\eta_{1})+\eta_{1}r^{2}}\frac{\gamma^{*}}{2}$. Moreover, when $\iota_{1}$ is even lower or equal to $\tfrac{r}{(1-\eta_{1})+\eta_{1}r^{2}}\frac{\gamma^{*}}{2}$, i.e. when condition ${\mathcal{R}e}(\gamma^{*}_{2})/\gamma^{*}<1/2$ does not hold, the power is asymptotically equal to 1 because the test statistic (3.4) diverges to infinity as a consequence of Remark 2.1.
3.1.2. Hypothesis test on $W$
As $W^{\top}{\boldsymbol{1}}={\boldsymbol{1}}$ (i.e. $w^{*}=1$ and $r=1$), we have
$w_{1,1}=1-w_{2,1}\quad\mbox{and}\quad w_{2,2}=1-w_{1,2}$
and so we can make inference on the two parameters $w_{2,1}$ and $w_{1,2}$. Using the parametration $\iota=w_{1,2}+w_{2,1}$ for the intensity of the interaction due to the matrix $W$, and $\eta=\tfrac{w_{1,2}}{w_{1,2}+w_{2,1}}$ (resp. $(1-\eta)$) for the percentage of this intensity due to the influence of process $1$ on process $2$ (resp., of process $2$ on process $1$), we have
(3.5) $w_{1,1}=1-(1-\eta)\iota,\qquad w_{1,2}=\eta\iota,\qquad w_{2,1}=(1-\eta)\iota\qquad\mbox{and}\qquad w_{2,2}=1-\eta\iota\,$,
where (by the required conditions on $W$)
(3.6) $\eta\in(0,1)\qquad\mbox{and}\qquad\iota\in\big{(}0,\min(1/\eta,1/(1-\eta)\big{)}\subseteq(0,2)\,.$
Two-sided test
If we take the null hypothesis $H_{0}:\;\iota=\iota_{0}\,,\,\eta=\eta_{0}$, where (3.6) and $\iota_{0}>1/2$ (the required condition in Assumption 2.2 for the second eigenvalue of $W$) are satisfied, then we can use Theorem 2.3 and obtain the test statistic (see (C.1) from Remark C.1 with $\phi=w^{*}=1$, $r=1$, $\phi_{2}^{*}=1-\iota_{0}$, g(x)=x(1x)g(x)=x(1-x) and ${\mathbf{B}}_{t}={\mathbf{K}}_{t}(c)/t$ with $c$ an arbitrary old color)
(3.7) $\Delta_{0}\,\frac{(K_{1,t}(c)-K_{2,t}(c))^{2}}{\widetilde{K}_{t}(c)(1-\widetilde{K}_{t}(c)/t)}\stackrel{{\scriptstyle d}}{{\longrightarrow}}\chi^{2}(1)\quad\mbox{under }H_{0}\,$,
$\mbox{where}\qquad\Delta_{0}=\frac{1}{2}-(1-\iota_{0})=\iota_{0}-\tfrac{1}{2},$
and
(3.8) $\widetilde{K}_{t}(c)={\mathbf{v}}^{\top}{\mathbf{K}}_{t}(c)=\eta_{0}K_{1,t}(c)+(1-\eta_{0})K_{2,t}(c).$
Since $K_{j,t}(c)/t\stackrel{{\scriptstyle a.s.}}{{\to}}\widetilde{P}_{\infty}(c)$, we obtain for any $\eta_{0}$ and $\eta_{1}$
$\frac{\eta_{1}K_{1,t}(c)/t+(1-\eta_{1})K_{2,t}(c)/t}{\widetilde{K}_{t}(c)/t}\,\frac{(1-\eta_{1}K_{1,t}(c)/t+(1-\eta_{1})K_{2,t}(c)/t)}{(1-\widetilde{K}_{t}(c)/t)}\stackrel{{\scriptstyle a.s.}}{{\longrightarrow}}1\,$,
and so the choice of the value of $\eta_{0}$ in $H_{0}$ does not affect the asymptotic distribution of the test statistic. This means that we can use the test statistic (3.7) with an arbitrary value of $\eta_{0}$ in order to perform a test on the intensity $\iota$. In particular, if we choose $\eta_{0}=1/2$, we simply have $\widetilde{K}_{t}(c)=\tfrac{K_{1,t}(c)+K_{2,t}(c)}{2}$.
One-sided test
The statistic (3.7) is increasing in $\iota_{0}$ and so it works well also for a one-sided test with the condition $\iota=\iota_{0}$ in $H_{0}$ replaced by $\iota\geq\iota_{0}$, where $\iota_{0}\in(\tfrac{1}{2},2)$.
Power
Following Remark B.1, we find that, under the alternative $H_{1}:\,\iota=\iota_{1}$, where $\iota_{1}\in(\tfrac{1}{2},2)$ with $\iota_{1}\neq\iota_{0}$, the statistic (3.7) is asymptotically distributed as
$\frac{\Delta_{0}}{\Delta_{1}}\chi^{2}(1)=\frac{\iota_{0}-\tfrac{1}{2}}{\iota_{1}-\tfrac{1}{2}}\,\chi^{2}(1)\,.$
Notice that the term $\tfrac{\Delta_{0}}{\Delta_{1}}$ is time-invariant and so, as before, increasing the time $t$ does not increase the power. However, the power is monotone decreasing in $\iota_{1}$, and it tends to $1$ as $\iota_{1}$ approaches its lower bound $1/2$, and it is equal to one for $\iota_{1}\leq 1/2$ by Remark 2.1.
3.1.3. Confidence interval for $\widetilde{P}_{\infty}(c)$
Fix the confidence level $(1-\alpha)$ and denote by $z_{\alpha}$ the quantile of order $(1-\alpha/2)$ of the standard normal distribution, that is ${\mathcal{N}}(0,1)(z_{\alpha},+\infty)=\alpha/2$. Assuming $\iota>1/2$, we can use Theorem 2.3 in order to construct a confidence interval for the random limit $\widetilde{P}_{\infty}(c)$ associated to a given item $c$ (see the interval (C.2) from Remark C.1 with $\phi=w^{*}=1$, $r=1$, g(x)=x(1x)g(x)=x(1-x) and ${\mathbf{B}}_{t}={\mathbf{K}}_{t}(c)/t$). Precisely, we obtain:
(3.9) $CI_{1-\alpha}(\widetilde{P}_{\infty}(c))\ =\ \frac{\widetilde{K}_{t}(c)}{t}\pm\ \frac{z_{\alpha}}{t^{1/2}}\!\sqrt{\frac{\widetilde{K}_{t}(c)}{t}\left(1-\frac{\widetilde{K}_{t}(c)}{t}\right)c_{\eta}},$
with $c_{\eta}=\eta^{2}+(1-\eta)^{2}$ and $\widetilde{K}_{t}(c)$ defined as in (3.8). It is interesting to notice the absence of the dependence on the specific value of $\iota$. On the other hand, while the dependence of $\widetilde{K}_{t}(c)$ on $\eta$ is not important, because $K_{j,t}(c)\stackrel{{\scriptstyle a.s.}}{{\to}}\widetilde{P}_{\infty}(c)$ so that $\frac{\eta_{1}K_{1,t}(c)/t+(1-\eta_{1})K_{2,t}(c)/t}{\widetilde{K}_{t}(c)/t}\stackrel{{\scriptstyle a.s.}}{{\to}}1$, the presence of the quantity $c_{\eta}$ implies that we need to know the value of the parameter $\eta$. However, $c_{\eta}$ is always smaller than $1$ for $\eta\in(0,1)$ and hence, if we do not know the value of $\eta$, we can use $\widetilde{K}_{t}(c)$ with $\eta=1/2$, so that we have $\widetilde{K}_{t}(c)=\tfrac{K_{1,t}(c)+K_{2,t}(c)}{2}$, and take the largest interval obtained from (3.9) with $c_{\eta}$ replaced by $1$.
As an example, we have simulated $S=200$ independent pair of innovation processes with $\Gamma$ and $W$ defined with $r=0.75$, $\gamma^{*}=0.75$, $\eta=1/2$, $\iota_{\Gamma}=1$ and $\iota_{W}=1.25$ and then, for any simulation, we have constructed the confidence intervals (3.9) of level $(1-\alpha)=95\%$ computed at time-step $t=10^{3}$ associated to the item $c$ with highest $(K_{t,1}(c)+K_{t,2}(c))$ (recall that the interval has the same expression for any given item $c$). In Figure 1 we can see these $S=200$ confidence intervals (sorted according to the value of $K_{t,1}(c)+K_{t,2}(c)\,$) and the estimates of the corresponding $\widetilde{P}_{\infty}(c)$ computed as $\tfrac{K_{t_{\infty},1}(c)+K_{t_{\infty},2}(c)}{2t_{\infty}}$ at time-step $t_{\infty}=10^{5}$. The empirical coverage is represented by the proportion of times the estimates of $\widetilde{P}_{\infty}(c)$ belong to the corresponing confidence interval, and in this simulation this is equal to $0.955$, which is very close to the nominal value of $0.95$.
Figure 1. Confidence interval (see (3.9)) for the random variable $\widetilde{P}_{\infty}(c)$ at time-step $t=10^{3}$ and related to item $c$ with highest ${\boldsymbol{1}}^{\top}{\mathbf{K}}_{t}(c)$ in the case $N=2$, $\Gamma$ defined as in (3.1) with $r=0.75$, $\gamma^{*}=0.75$, $\eta=1/2$ and $\iota_{\Gamma}=1$, and $W$ defined as in (3.5) with $\eta=1/2$ and $\iota_{W}=1.25$. Number of simulations $S=200$ (sorted according to the value of ${\boldsymbol{1}}^{\top}{\mathbf{K}}_{t}(c)$). The crosses represent the value of $\widetilde{P}_{\infty}(c)$ estimated by ${\boldsymbol{1}}^{\top}{\mathbf{K}}_{t_{\infty}}(c)/(Nt_{\infty})$ at time-step $t_{\infty}=10^{5}$.
3.2. Mean-field interaction for $N\geq 2$
3.2.1. Hypothesis test on $\Gamma$
Suppose that $\Gamma$ is of the form
(3.10) $\gamma_{j,h}=\phi\left(\iota/N+\delta_{j,h}(1-\iota)\right)\qquad\phi,\iota\in(0,1]\,.$
We refer to this kind of interaction matrix as the mean-field interaction because in this case the term $\sum_{j=1}^{N}\gamma_{j,h}D^{*}_{t,j}$ in (1.3) becomes $\phi\left(\iota\frac{1}{N}\sum_{j=1}^{N}D^{*}_{t,j}+(1-\iota)D^{*}_{t,h}\right)$ and so the conditional probability $Z^{*}_{t,h}$ depends on a convex combination between $D^{*}_{t,h}$ and the averaged value of the all $D^{*}_{t,j}$ in the system. Note that, for this matrix $\Gamma$, the leading eigenvalue $\gamma^{*}$ coincides with the parameter $\phi$.
We can make inference on the parameter $\iota$ that rules the intensity of the interaction.
Two-sided test
If we take the null hypothesis $H_{0}:\;\iota=\iota_{0}$, where $\iota_{0}\leq 1$ and $\iota_{0}>1/2$ (the required condition in Assumption 2.1 for the second eigenvalue of $\Gamma$), then we can use Theorem 2.1 and obtain the test statistic (see (C.4) from Remark C.2 with $\phi=\gamma^{*}$, $\phi_{2}^{*}=\gamma^{*}(1-\iota_{0})$, $g(x)=x$, and ${\mathbf{B}}_{t}=\frac{{\mathbf{D}}_{t}^{*}}{\zeta_{t}(\gamma^{*})}$ or, equivalently ${\mathbf{B}}_{t}=\frac{{\mathbf{D}}_{t}^{*}}{t^{\gamma^{*}}}$)
(3.11) $2\Delta_{0}\frac{\big{\|}{\mathbf{D}}^{*}_{t}-\widetilde{D}^{*}_{t}{\boldsymbol{1}}\big{\|}^{2}}{\widetilde{D}^{*}_{t}}\stackrel{{\scriptstyle d}}{{\longrightarrow}}\chi^{2}(N-1)\quad\mbox{under }H_{0},$
$\mbox{where}\qquad\Delta_{0}=\frac{1}{2}-\frac{\gamma^{*}(1-\iota_{0})}{\gamma^{*}}=\iota_{0}-\frac{1}{2}$
and $\widetilde{D}^{*}_{t}={\mathbf{v}}^{\top}{\mathbf{D}}^{*}_{t}=\tfrac{{\boldsymbol{1}}^{\top}{\mathbf{D}}^{*}_{t}}{N}$. Note that this test statistic does not depend on the parameter $\phi=\gamma^{*}$ and so the value of this parameter is not needed for the test on $\iota$.
One-sided test
Since the test statistic (3.11) is increasing in $\iota_{0}$, it works well also for the one-sided test with $H_{0}:\;\iota\geq\iota_{0}$, where $\iota_{0}\in(1/2,1]$.
Power
Following Remark B.1, we observe that, under the alternative $H_{1}:\,\iota=\iota_{1}$, with $\iota_{1}\in(1/2,1]$, the statistics (3.11) is asymptotically distributed as
$\frac{\Delta_{0}}{\Delta_{1}}\chi^{2}(N-1)=\frac{\iota_{0}-\tfrac{1}{2}}{\iota_{1}-\tfrac{1}{2}}\chi^{2}(N-1)\,.$
Therefore, as before, increasing the time $t$ does not increase the power. Moreover, the power is monotone decreasing in $\iota_{1}$, and it tends to $1$ as $\iota_{1}$ approaches its lower bound $1/2$, being equal to 1 for $\iota_{1}\leq 1/2$ by Remark 2.1.
3.2.2. Hypothesis test on $W$
Suppose that $W$ is of the form
(3.12) $w_{j,h}=\iota/N+\delta_{j,h}(1-\iota)\qquad\iota\in(0,1]\,.$
Two-sided test
If we take the null hypothesis $H_{0}:\;\iota=\iota_{0}$, where $\iota_{0}\leq 1$ and $\iota_{0}>1/2$ (the required condition in Assumption 2.2 for the second eigenvalue of $W$), then we can use Theorem 2.3 and obtain the test statistic (see (C.4) from Remark C.2 with $\phi=w^{*}=1$, $\phi_{2}^{*}=w^{*}(1-\iota_{0})=(1-\iota_{0})$, g(x)=x(1x)g(x)=x(1-x), and ${\mathbf{B}}_{t}={\mathbf{K}}_{t}(c)/t$)
(3.13) $2\Delta_{0}\frac{\big{\|}{\mathbf{K}}_{t}(c)-\widetilde{K}_{t}(c){\boldsymbol{1}}\big{\|}^{2}}{\widetilde{K}_{t}(c)(1-\widetilde{K}_{t}(c)/t)}\stackrel{{\scriptstyle d}}{{\longrightarrow}}\chi^{2}(N-1)\quad\mbox{under }H_{0},$
$\mbox{where}\qquad\Delta_{0}=\frac{1}{2}-(1-\iota_{0})=\iota_{0}-\frac{1}{2},$
$c$ is an arbitrary observed item and $\widetilde{K}_{t}(c)={\mathbf{v}}^{\top}{\mathbf{K}}_{t}(c)=\tfrac{{\boldsymbol{1}}^{\top}{\mathbf{K}}_{t}(c)}{N}$.
One-sided test
The test statistic (3.13) is increasing in $\iota_{0}$ and so it works well also for the one-sided test with $H_{0}:\;\iota\geq\iota_{0}$, where $\iota_{0}\in(1/2,1]$.
Power
Following Remark B.1, we observe that, under the alternative $H_{1}:\,\iota=\iota_{1}$, with $\iota_{1}\in(1/2,1]$, the statistics (3.13) is asymptotically distributed exactly as the previous statistics (3.11) and so the same considerations on the power of the test hold true.
3.2.3. Confidence interval for $\widetilde{P}_{\infty}(c)$
Fix the confidence level $(1-\alpha)$ and denote by $z_{\alpha}$ the quantile of order $(1-\alpha/2)$ of the standard normal distribution, that is ${\mathcal{N}}(0,1)(z_{\alpha},+\infty)=\alpha/2$. Assuming $\iota\in(1/2,1]$, we can use Theorem 2.3 in order to construct a confidence interval for the random limit $\widetilde{P}_{\infty}(c)$ associated to a given item $c$ (see the interval (C.3) from Remark C.2 with $\phi=w^{*}=1$, g(x)=x(1x)g(x)=x(1-x) and ${\mathbf{B}}_{t}={\mathbf{K}}_{t}(c)/t$). Precisely, we obtain:
(3.14) $CI_{1-\alpha}(\widetilde{P}_{\infty}(c))\ =\ \frac{\widetilde{K}_{t}(c)}{t}\pm\ \frac{z_{\alpha}}{t^{1/2}}\!\sqrt{\frac{\widetilde{K}_{t}(c)}{t}\left(1-\frac{\widetilde{K}_{t}(c)}{t}\right)\frac{1}{N}},$
where $\widetilde{K}_{t}(c)={\mathbf{v}}^{\top}{\mathbf{K}}_{t}(c)=\tfrac{{\boldsymbol{1}}^{\top}{\mathbf{K}}_{t}(c)}{N}$. Note that this interval does not depend on the value of the parameter $\iota$.
As an example, we have simulated $S=200$ independent triplets of innovation processes with $\Gamma$ and $W$ of the mean-field type defined with $N=3$, $\gamma^{*}=0.75$, $\iota_{\Gamma}=\iota_{W}=0.8$, and we show the confidence intervals (3.14) of level $(1-\alpha)=95\%$ at time-step $t=10^{3}$, associated to the item $c$ with highest ${\boldsymbol{1}}^{\top}{\mathbf{K}}_{t}(c)$ (see Figure 2). In this case, the empirical coverage is equal to $0.970$, which is again very close to the nominal value $0.95$.
Figure 2. Confidence interval (see (3.14)) for the random variable $\widetilde{P}_{\infty}(c)$ at time-step $t=10^{3}$ and related to item $c$ with highest ${\boldsymbol{1}}^{\top}{\mathbf{K}}_{t}(c)$ in the case $N=3$, $\Gamma$ defined as in (3.10) with $\phi=0.75$ and $\iota_{\Gamma}=0.8$, and $W$ defined as in (3.12) with $\iota_{W}=0.8$. Number of simulations $S=200$ (sorted according to the value of ${\boldsymbol{1}}^{\top}{\mathbf{K}}_{t}(c)$). The crosses represent the value of $\widetilde{P}_{\infty}(c)$ estimated by ${\boldsymbol{1}}^{\top}{\mathbf{K}}_{t_{\infty}}(c)/(Nt_{\infty})$ at time-step $t_{\infty}=10^{5}$.
4. Applications
In this section we apply the proposed inferential methodology on the two real data sets, already considered in Aletti et al. (2023), to test the minimum level of interaction presented by the matrices $\Gamma$ and $W$ in these specific contexts. In particular, the first one is taken from the social content aggregation website Reddit, collected, elaborated and made freely available on the web by the authors of Monti et al. (2023) (see https://github.com/corradomonti/demographic-homophily), while the second one is downloaded from the on-line library Project Gutenberg, which is a collection of public domain books (see https://www.gutenberg.org).
4.1. Reddit data set
This data set consists of a collection of news, and comments associated to each news, for the period 201620202016-2020, downloaded from the r/news community on the website Reddit Monti et al. (2023). Each news is associated with the author who posted it. Moreover, the data set contains the specific topic the news belongs to and, to each comment is also assigned a measurement of the sentiment, expressed as a real value in (1,1)(-1,1). We consider only the comments to news belonging to the topic ”Politics“. Moreover, we categorize the sentiment variable: precisely, we define it as ”positive“ if the provided sentiment value was larger than +0.35+0.35 and ”negative“ if the provided sentiment value was lower than 0.35-0.35. (We remove any comment with an original sentiment value that lies within (0.35,+0.35)(-0.35,+0.35).) We finally take a data set of 2 602 1732\,602\,173 comments for each of the two categories. Hence, we consider two innovation processes: one related to the authors who receive at least one comment with negative sentiment for the news they post and the similar one referred to the comments with positive sentiment. A first analysis of this data set can be found in Aletti et al. (2023), where we illustrate that these two sequences exhibit empirical behaviours that are in accordance with those predicted by Theorems 1.1 and 1.2, and we provide the estimates $\widehat{\gamma}^{*}=0.781$ and $\widehat{r}=10^{-0.727}=0.187$ (these estimates have been obtained by means of a suitable regression on $({\mathbf{D}}^{*}_{t})_{t}$, which simulations have shown to be a highly accurate estimation method). We here continue the study performing the two hypothesis tests on $\Gamma$ (taking the above estimates as the known values of $\gamma^{*}$ and $r$) and on $W$ described in Subsection 3.1. More precisely, we consider the two one-sided tests:
$H_{\Gamma,0}:\;\iota\geq\iota_{\Gamma,0},\,\eta\leq 1/2\qquad\mbox{and}\qquad H_{W,0}:\;\iota\geq\iota_{W,0},$
where $\iota_{\Gamma,0}$ must be selected in the interval (0.141, 0.291](0.141,\,0.291] and $\iota_{W,0}$ must be selected in the interval $(\tfrac{1}{2},\,2)$. The $p$-values associated to the null hypothesis $H_{\Gamma,0}$ and $H_{W,0}$, for different values of $\iota_{\Gamma,0}$ and $\iota_{W,0}$, respectively, are collected in Table 1. In order to better describe the empirical distribution of the $p$-values for testing $W$, instead of taking a single item $c$, we take the $100$ items $c$ with highest $(K_{t,1}(c)+K_{t,2}(c))$, and report in Table 1 the summary statistics of this sample (i.e. the quantiles of order $50\%$, $75\%$ and $95\%$). We can see that, considering a test on $\Gamma$ performed at level $\alpha=0.05$, we can reject the null hypothesis with $\iota_{\Gamma,0}=0.21$, but we cannot reject it for $\iota_{\Gamma,0}=0.18$. Regarding the test on $W$ instead, we observe that the proportion of items $c$ providing a $p$-value that let us reject the null hypothesis that the interaction intensity $\iota_{W}$ is higher than $\iota_{W,0}$ is from $75\%$ to $95\%$ for $\iota_{W,0}\leq 1.25$, while is even higher than $95\%$ or $\iota_{W,0}=1.50$.
Table 1. Reddit data set. $p$-values associated to the hypothesis tests described in Subsection 3.1 with $H_{\Gamma,0}:\iota_{\Gamma}\geq\iota_{\Gamma,0},\,\eta\leq 1/2$ and $H_{W,0}:\iota_{W}\geq\iota_{W,0}$, respectively. The table also contains the summary statistics related to the sample of the $100$ items $c$ with highest $(K_{t,1}(c)+K_{t,2}(c))$.
$\iota_{\Gamma,0}$ p-value $\iota_{W,0}$ p-value: $q_{50\%}$,  $q_{75\%}$,  $q_{95\%}$
0.15 0.384 0.60 <<0.001, 0.007, 0.486
0.18 0.064 0.80 <<0.001, <<0.001, 0.229
0.21 0.014 1.00 <<0.001, <<0.001, 0.121
0.24 0.003 1.25 <<0.001, <<0.001, 0.058
0.27 0.001 1.50 <<0.001, <<0.001, 0.029
4.2. Gutenberg data set
This data set comes from the on-line library Project Gutenberg and it is related to seven books written in English and belonging to two particular literary genres: “Western” and “’History”. Precisely, it consists of 476 948476\,948 words for both genres. The two sequences of words are the innovation processes object of the study. Therefore, each word is new or old for a specific literary genre if, respectively, it has or has not already been used within that genre. Analogously, each word is new or old for the entire system if, respectively, it has or has not already been used within any considered genre. A first analysis of also this data set can be found in Aletti et al. (2023), where we show that the behaviors of these two sequences predicted by Theorems 1.1 and 1.2 match with the observed ones, and we provide the estimates $\widehat{\gamma}^{*}=0.466$ and $\widehat{r}=10^{-0.238}=0.578$. As in the previous application, we here continue the study performing the two hypothesis tests on $\Gamma$ and $W$ described in Subsection 3.1. More precisely, we consider the two one-sided tests:
$H_{\Gamma,0}:\;\iota\geq\iota_{0},\,\eta\leq 1/2\qquad\mbox{and}\qquad H_{W,0}:\;\iota\geq\iota_{0},$
where $\iota_{\Gamma,0}$ must be selected in the interval (0.202, 0.539](0.202,\,0.539] and $\iota_{W,0}$ must be selected in the interval $(\tfrac{1}{2},\,2)$. The $p$-values associated to the null hypothesis $H_{\Gamma,0}$ and $H_{W,0}$, for different values of $\iota_{\Gamma,0}$ and $\iota_{W,0}$, respectively, are collected in Table 2. Moreover, we take the $100$ items $c$ with highest $(K_{t,1}(c)+K_{t,2}(c))$ and report in the table the empirical quantiles of the corresponding sample of $p$-values. We can observe that, considering a test on $\Gamma$ performed at level $\alpha=0.05$, we can never reject the null hypothesis. Regarding the test on $W$ instead, we observe that the proportion of items $c$ providing a $p$-value that let us reject the null hypothesis that the interaction intensity $\iota_{W}$ is higher than $\iota_{W,0}$ is from $50\%$ to $75\%$ for $\iota_{W,0}=0.60$, while is from $75\%$ to $95\%$ for $\iota_{W,0}\geq 0.80$.
Table 2. Gutenberg data set. $p$-values associated to the hypothesis tests described in Subsection 3.1 with $H_{\Gamma,0}:\iota_{\Gamma}\geq\iota_{\Gamma,0},\,\eta\leq 1/2$ and $H_{W,0}:\iota_{W}\geq\iota_{W,0}$, respectively. The table also contains the summary statistics related to the sample of the $100$ items $c$ with highest $(K_{t,1}(c)+K_{t,2}(c))$.
$\iota_{\Gamma,0}$ p-value $\iota_{W,0}$ p-value: $q_{50\%}$,  $q_{75\%}$,  $q_{95\%}$
0.21 0.034 0.60 0.010, 0.209, 0.823
0.24 <<0.001 0.80 <<0.001, 0.030, 0.699
0.27 <<0.001 1.00 <<0.001, 0.005, 0.618
0.30 <<0.001 1.25 <<0.001, 0.001, 0.541
0.40 <<0.001 1.50 <<0.001, <<0.001, 0.480
Author contributions statement
All the authors contributed equally to the present work.
Notation
In all the appendix, the notation $R_{t}=O(s_{t})$ means that $|R_{t}|\leq Cs_{t}$ for a suitable constant $C$ and for $t$ large enough. Similarly, the vectorial notation ${\mathbf{R}}_{t}=O(s_{t}){\boldsymbol{1}}$ means that each component of ${\mathbf{R}}_{t}$, say $R_{t,j}$, is such that $|R_{t,j}|\leq Cs_{t}$ for a suitable constant $C$ and for $t$ large enough.
Appendix A General results
Consider two multi-dimensional real stochastic processes $({\mathbf{A}}_{t},{\mathbf{B}}_{t})_{t}$, adapted to a filtration $(\mathcal{F}_{t})_{t}$, with the following joint dynamics:
[…]
where ${\mathbf{A}}_{0}$ and ${\mathbf{B}}_{0}$ are integrable and
  • $(i)$
    $\Phi^{\top}$ is a non-negative irreducible matrix with leading eigenvalue $0<\phi^{*}\leq 1$;
  • $(ii)$
    ${\mathbf{R}}_{A,t+1}=O(t^{-(1+\beta)}){\boldsymbol{1}}$ and ${\mathbf{R}}_{B,t+1}=O(t^{-(1+\beta)}){\boldsymbol{1}}$ for some $\beta>\frac{\phi^{*}}{2}$;
  • $(iii)$
    $\Delta{\mathbf{M}}_{t+1}=O(t^{1-\eta}){\boldsymbol{1}}$ with $\eta>\phi^{*}/2$ and $t^{-(1-\phi^{*})}E[\Delta{{\mathbf{M}}}_{t+1}\Delta{{\mathbf{M}}}_{t+1}^{\top}|\mathcal{F}_{t}]\stackrel{{\scriptstyle a.s.}}{{\rightarrow}}\Sigma_{\infty}$ with $\Sigma_{\infty}$ a positive semidefinite symmetric random matrix.
It is worthwhile to note that dynamics similar to (A.1) have already been considered in the Stochastic Approximation literature (e.g. Laruelle and Pagès (2013); Mokkadem and Pelletier (2006)). However, this kind of results do not apply to our situation. Indeed, the covariance matrices in those results are deterministic, while the asymptotic covariance matrices in our CLTs are random. This is also why we do not use the simple convergence in distribution, but we employ the notion of stable convergence, which is, among other things, essential for obtaining the statistical tools.
Let $\mathbf{u}$} and $\mathbf{v}$} be the right and the left eigenvectors of $\Phi^{\top}$ associated to $\phi^{*}$ and such that ${\mathbf{v}}^{\top}{\boldsymbol{1}}=1\quad\mbox{and}\quad{\mathbf{v}}^{\top}{\mathbf{u}}=1.$ (Recall that the components of these vectors are all strictly positive according to the Frobenius-Perron theory.)
Let $\widetilde{A}_{t}={\mathbf{v}}^{\top}{\mathbf{A}}_{t}$ and get its dynamics by multypling $\mathbf{v}$$^{\top}$ to the dynamics of ${\mathbf{A}}_{t}$ as follows:
$\widetilde{A}_{t+1}=\widetilde{A}_{t}+\frac{1}{t+1}\phi^{*}{\mathbf{v}}^{\top}\Delta{{\mathbf{M}}}_{t+1}+{\mathbf{v}}^{\top}{{\mathbf{R}}}_{A,t+1}.$
Define
(A.2) $\widehat{{\mathbf{A}}}_{t}={\mathbf{A}}_{t}-\widetilde{A}_{t}{\mathbf{u}}=(I-{\mathbf{u}}{\mathbf{v}}^{\top}){\mathbf{A}}_{t}\qquad\mbox{and}\qquad\widehat{{\mathbf{B}}}_{t}={\mathbf{B}}_{t}-\frac{1}{\phi^{*}}\widetilde{A}_{t}{\mathbf{u}}.$
Theorem A.1.
(CLT for $(\widetilde{A}_{t})$ when $N\geq 1$)
In addition to assumptions $(i)$, $(ii)$ and $(iii)$, also assume:
  • $(iv)$
    $\widetilde{A}_{t}\stackrel{{\scriptstyle a.s./in\,mean}}{{\longrightarrow}}\widetilde{A}_{\infty}$, where $\widetilde{A}_{\infty}$ is an integrable random variable.
Then, we have
$t^{\phi^{*}/2}(\widetilde{A}_{t}-\widetilde{A}_{\infty})\stackrel{{\scriptstyle stably}}{{\longrightarrow}}\mathcal{N}(0,\phi^{*}{\mathbf{v}}^{\top}\Sigma_{\infty}{\mathbf{v}})\,.$
Moreover, the above convergence is in the sense of the almost sure conditional convergence with respect to the filtration $(\mathcal{F}_{t})$.
Proof.
Set $\breve{M}_{t}=\sum_{n=1}^{t}\frac{1}{n}{\mathbf{v}}^{\top}\Delta{{\mathbf{M}}}_{n}$ and $\widetilde{R}_{A,t+1}={\mathbf{v}}^{\top}{{\mathbf{R}}}_{A,t+1}$. Therefore, we have
$\begin{split}\widetilde{A}_{t+1}-\widetilde{A}_{0}&=\sum_{n=0}^{t}\widetilde{A}_{n+1}-\widetilde{A}_{n}\\ &=\phi^{*}\sum_{n=0}^{t}\frac{1}{n+1}{\mathbf{v}}^{\top}\Delta{{\mathbf{M}}}_{n+1}+\sum_{n=0}^{t}\widetilde{R}_{A,n+1}\\ &=\phi^{*}\breve{M}_{t+1}+\sum_{n=0}^{t}\widetilde{R}_{A,n+1}\,.\end{split}$
Since $\widetilde{A}_{t}\stackrel{{\scriptstyle a.s./in\,mean}}{{\to}}\widetilde{A}_{\infty}$ and $\sum_{n\geq 0}\widetilde{R}_{A,n+1}=\sum_{n\geq 0}O(1/n^{1+\beta})<+\infty$  then $\breve{M}_{t}\stackrel{{\scriptstyle a.s./in\,mean}}{{\to}}\breve{M}_{\infty}$ for a certain integrable random variable $\breve{M}_{\infty}$. Moreover, we have
[…]
where $\beta-\phi^{*}/2>0$. Summing up it is enough to prove the convergence of $t^{\phi^{*}/2}(\breve{M}_{t}-\breve{M}_{\infty})$. For this purpose, we want to apply Theorem D.1. Let us consider, for each $t$ the filtration $({\mathcal{F}}_{t,k})_{k}$ and the process $(L_{t,k})_{k}$ defined by
${\mathcal{F}}_{t,0}={\mathcal{F}}_{t,1}={\mathcal{F}}_{t},\qquad L_{t,0}=L_{t,1}=0$
and, for $k\geq 2$,
${\mathcal{F}}_{t,k}={\mathcal{F}}_{t+k-1},\qquad L_{t,k}=t^{\phi^{*}/2}(\breve{M}_{t}-\breve{M}_{t+k-1}).$
The process $(L_{t,k})_{k}$ is a martingale w.r.t. $({\mathcal{F}}_{t,k})_{k}$ which converges (for $k\to+\infty$) a.s. and in mean to the random variable $L_{t,\infty}=t^{\phi^{*}/2}(\breve{M}_{t}-\breve{M}_{\infty})$. In addition, the increment $Y_{t,n}=L_{t,n}-L_{t,n-1}$ is equal to zero for $n=1$ and, for $n\geq 2$, it coincides with a random variable of the form $t^{\phi^{*}/2}(\breve{M}_{k}-\breve{M}_{k+1})$ with $k\geq t$. Therefore, we have
[…]
where the almost sure convergence follows by applying (Crimaldi et al., 2016, Lemma 4.1) and by the fact that
$k^{-(1-\phi^{*})}E[({\mathbf{v}}^{\top}\Delta{{\mathbf{M}}}_{k+1})^{2}|\mathcal{F}_{t}]=k^{-(1-\phi^{*})}{\mathbf{v}}^{\top}E[\Delta{{\mathbf{M}}}_{k+1}\Delta{\mathbf{M}}_{k+1}^{\top}|\mathcal{F}_{t}]{\mathbf{v}}\stackrel{{\scriptstyle a.s.}}{{\longrightarrow}}{\mathbf{v}}^{\top}\Sigma_{\infty}{\mathbf{v}}\,.$
Finally, we have
$\sup_{n\geq 1}|Y_{t,n}|=t^{\phi^{*}/2}\,\sup_{k\geq t}|\breve{M}_{k}-\breve{M}_{k+1}|\leq\sup_{k\geq t}\,k^{\phi^{*}/2}\frac{|{\mathbf{v}}^{\top}\Delta{\mathbf{M}}_{k+1}|}{(k+1)}$
and, since $\Delta{\mathbf{M}}_{k+1}=O((k+1)^{1-\eta}){\boldsymbol{1}}$ by assumption $(iii)$ and so $k^{\phi^{*}/2}\frac{|{\mathbf{v}}^{\top}\Delta{\mathbf{M}}_{k+1}|}{(k+1)}=O(k^{-(\eta-\phi^{*}/2)})$, the thesis follows by Theorem D.1. ∎
Theorem A.2.
(Convergence result for $(\widehat{A}_{t},\widehat{B}_{t})$ when $N=1$)
Under the assumptions $(i)$ and $(ii)$, we have
$\widehat{A}_{t}\equiv 0\qquad\mbox{and}\qquad\widehat{B}_{t}=B_{t}-\frac{1}{\phi^{*}}A_{t}\stackrel{{\scriptstyle a.s.}}{{=}}o(t^{-\frac{\phi^{*}}{2}}).$
Proof.
First, when $N=1$ it is immediate to see $\widetilde{A}_{t}\equiv A_{t}$ and hence $\widehat{A}_{t}=A_{t}-\widetilde{A}_{t}\equiv 0$ for any $t\geq 1$. The dynamics of $\widehat{B}_{t}$ can be obtained by subtracting the dynamics of $\frac{1}{\phi^{*}}\widetilde{A}_{t}=\frac{1}{\phi^{*}}A_{t}$ to the dynamics of $B_{t}$ as follows:
$\widehat{B}_{t+1}=\widehat{B}_{t}-\frac{1}{t+1}(\phi^{*}\widehat{B}_{t}+\phi^{*}\frac{1}{\phi^{*}}\widetilde{A}_{t}-A_{t})+\frac{1}{t+1}\Delta{M}_{t+1}-\frac{1}{\phi^{*}}(\widetilde{A}_{t+1}-\widetilde{A}_{t})+{R}_{B,t+1}\,$,
from which we can derive
$\widehat{B}_{t+1}=\widehat{B}_{t}-\frac{1}{t+1}(\phi^{*}\widehat{B}_{t})+\frac{1}{t+1}\Delta{M}_{t+1}-\frac{1}{\phi^{*}}\left(\frac{1}{t+1}\phi^{*}\Delta{M}_{t+1}\right)+{R}_{B,t+1}-\frac{1}{\phi^{*}}{R}_{A,t+1}\,$,
that is
$\widehat{B}_{t+1}=\left(1-\frac{\phi^{*}}{t+1}\right)\widehat{B}_{t}+O(t^{-(1+\beta)}).$
Then, since $\sum_{t}\frac{\phi^{*}}{t+1}=+\infty$ and $\sum_{t}O(t^{-(1+\beta)})<+\infty$, we can apply Lemma D.2 to obtain $\widehat{B}_{t}\stackrel{{\scriptstyle a.s.}}{{\rightarrow}}0$. Moreover, if we multiply the previous dynamics of $\widehat{B}_{t}$ by $(t+1)^{\frac{\phi^{*}}{2}}$ and we use the relation $(\frac{t+1}{t})^{\frac{\phi^{*}}{2}}=(1+\frac{1}{t})^{\frac{\phi^{*}}{2}}=1+\frac{\phi^{*}/2}{t+1}+O(1/t^{2})$, we get the following dynamics for the process $\widehat{B}^{*}_{t}=t^{\frac{\phi^{*}}{2}}\widehat{B}_{t}$:
$\widehat{B}^{*}_{t+1}=\left(1-\frac{\phi^{*}}{t+1}\right)\left(1+\frac{\phi^{*}/2}{t+1}\right)\widehat{B}_{t}^{*}+O(t^{-(1+\beta-\frac{\phi^{*}}{2})})+O(t^{-(2-\frac{\phi^{*}}{2})})\,$,
which is equivalent to
$\widehat{B}^{*}_{t+1}=\left(1-\frac{\phi^{*}/2}{t+1}\right)\widehat{B}_{t}^{*}+O(t^{-(1+\beta-\frac{\phi^{*}}{2})})+O(t^{-(2-\frac{\phi^{*}}{2})}).$
Therefore, since $\sum_{t}t^{-(2-\frac{\phi^{*}}{2})}<+\infty$ because $\phi^{*}/2<1$ and $\sum_{t}t^{-(1+\beta-\frac{\phi^{*}}{2})}<+\infty$ because $\beta>\phi^{*}/2$ by assumption, we can apply Lemma D.2 also to the sequence $\widehat{B}^{*}_{t}$ to obtain $\widehat{B}^{*}_{t}\stackrel{{\scriptstyle a.s.}}{{\rightarrow}}0$, i.e. $\widehat{B}_{t}\stackrel{{\scriptstyle a.s.}}{{=}}o(t^{-\frac{\phi^{*}}{2}})$. ∎
Theorem A.3.
(CLT for $(\widehat{{\mathbf{A}}}_{t},\widehat{{\mathbf{B}}}_{t})$ when $N\geq 2$)
In addition to assumptions $(i)$, $(ii)$ and $(iii)$, let $\phi_{2}^{*}$ be an eigenvalue of $\Phi$ different from $\phi^{*}$ with highest real part, i.e. $\phi_{2}^{*}\in Sp(\Phi)\setminus\{\phi^{*}\}$ such that $\mathcal{R}e(\phi_{2}^{*})=\max\{\mathcal{R}e(\phi):\,\phi\in Sp(\Phi)\setminus\{\phi^{*}\}\}$ and assume
  • $(v)$
    ${\mathcal{R}e}(\phi_{2}^{*})/\phi^{*}<1/{2}$.
Then, we have
$t^{\phi^{*}/2}(\widehat{{\mathbf{A}}}_{t},\widehat{{\mathbf{B}}}_{t})\stackrel{{\scriptstyle stably}}{{\longrightarrow}}\mathcal{N}({\boldsymbol{0}},{\mathcal{M}}_{\infty})$
with
(A.3) $\mathcal{M}_{\infty}=\begin{pmatrix}{\mathcal{M}}_{\infty}^{11}&{\mathcal{M}}_{\infty}^{13}\\ {\mathcal{M}}_{\infty}^{13}&{\mathcal{M}}_{\infty}^{33}\end{pmatrix}=\begin{pmatrix}U{\mathcal{S}}_{\infty}^{11}U^{\top}&U{\mathcal{S}}_{\infty}^{13}U^{\top}\\ U{\mathcal{S}}_{\infty}^{13}U^{\top}&U{\mathcal{S}}_{\infty}^{33}U^{\top}\end{pmatrix}\,$,
where, denoting by $\phi_{j}$ the eigenvalues of $\Phi$ different from $\phi^{*}$, we have for $j=2,\dots,N$
$\begin{split}[{\mathcal{S}}_{\infty}^{11}]_{h,j}&=\frac{\phi_{j}\phi_{h}}{\phi^{*}-\phi_{j}-\phi_{h}}({\mathbf{v}}_{j}^{\top}\Sigma_{\infty}{\mathbf{v}}_{h}),\\ [{\mathcal{S}}_{\infty}^{33}]_{h,j}&=\frac{1}{\phi^{*}-\phi_{j}-\phi_{h}}({\mathbf{v}}_{j}^{\top}\Sigma_{\infty}{\mathbf{v}}_{h}),\\ [{\mathcal{S}}_{\infty}^{13}]_{h,j}&=\frac{\phi_{j}}{\phi^{*}-\phi_{j}-\phi_{h}}({\mathbf{v}}_{j}^{\top}\Sigma_{\infty}{\mathbf{v}}_{h}).\end{split}$
Proof.
We split the proofs in some steps.
Step 1 (Dynamics of $\theta_{t}=(\widehat{A}_{t},\widehat{B}_{t})$).
Firstly, we recall that the matrix $\Phi^{\top}$ can be decomposed as
$\Phi^{\top}=\phi^{*}{\mathbf{u}}{\mathbf{v}}^{\top}+UDV^{\top}\,$,
where $D$ is the diagonal matrix whose elements are the eigenvalues of $\Phi$ (i.e. of $\Phi^{\top}$) different from $\phi^{*}$ and $U$ and $V$ denote the matrices whose columns are the left (right) and the right (left) eigenvectors of $\Phi$ (of $\Phi^{\top}$, respectively) associated to these eigenvalues, so that we have
(A.4) $V^{\top}{\mathbf{u}}=U^{\top}{\mathbf{v}}=0,\quad V^{\top}U=U^{\top}V=I,\quad I={\mathbf{u}}{\mathbf{v}}^{\top}+UV^{\top}\,.$
Then, we have $\widehat{{\mathbf{A}}}_{t}=UV^{\top}{\mathbf{A}}_{t}$ and the dynamics of $\widehat{{\mathbf{A}}}_{t}$ can be obtained by multypling by $UV^{\top}$ the dynamics of ${\mathbf{A}}_{t}$ as follows:
$\widehat{{\mathbf{A}}}_{t+1}=\widehat{{\mathbf{A}}}_{t}-\frac{1}{t+1}U(I\phi^{*}-D)V^{\top}\widehat{{\mathbf{A}}}_{t}+\frac{1}{t+1}UDV^{\top}\Delta{{\mathbf{M}}}_{t+1}+UV^{\top}{{\mathbf{R}}}_{A,t+1}.$
The dynamics of $\widehat{{\mathbf{B}}}_{t}$ can be obtained by subtracting the dynamics of $\frac{1}{\phi^{*}}\widetilde{A}_{t}{\mathbf{u}}$ to the dynamics of ${\mathbf{B}}_{t}$ as follows:
$\widehat{{\mathbf{B}}}_{t+1}=\widehat{{\mathbf{B}}}_{t}-\frac{1}{t+1}(\phi^{*}\widehat{{\mathbf{B}}}_{t}+\phi^{*}\frac{1}{\phi^{*}}\widetilde{A}_{t}{\mathbf{u}}-\widehat{{\mathbf{A}}}_{t}-\widetilde{A}_{t}{\mathbf{u}})+\frac{1}{t+1}\Delta{{\mathbf{M}}}_{t+1}-\frac{1}{\phi^{*}}{\mathbf{u}}(\widetilde{A}_{t+1}-\widetilde{A}_{t})+{{\mathbf{R}}}_{B,t+1}\,$,
from which we can derive
$\widehat{{\mathbf{B}}}_{t+1}=\widehat{{\mathbf{B}}}_{t}-\frac{1}{t+1}(\phi^{*}\widehat{{\mathbf{B}}}_{t}-\widehat{{\mathbf{A}}}_{t})+\frac{1}{t+1}(I-{\mathbf{u}}{\mathbf{v}}^{\top})\Delta{{\mathbf{M}}}_{t+1}+{{\mathbf{R}}}_{B,t+1}-\frac{1}{\phi^{*}}{\mathbf{u}}{\mathbf{v}}^{\top}{{\mathbf{R}}}_{A,t+1}\,.$
Then the dynamics of ${\boldsymbol{\eta}}_{t}=(\widehat{{\mathbf{A}}}_{t},\widehat{{\mathbf{B}}}_{t})$ can be expressed as
${\boldsymbol{\eta}}_{t+1}-{\boldsymbol{\eta}}_{t}=-\frac{1}{t+1}Q{\boldsymbol{\eta}}_{t}+\frac{1}{t+1}R\Delta{\mathbf{M}}_{{\boldsymbol{\eta}},t+1}+{\mathbf{R}}_{{\boldsymbol{\eta}},t+1}\,$,
where [^3] The symbol $\mathcal{O}$ denotes the null matrix.
$Q=\begin{pmatrix}U(I\phi^{*}-D)V^{\top}&\mathcal{O}\\ -UV^{\top}&I\phi^{*}\end{pmatrix},\qquad\Delta{\mathbf{M}}_{{\boldsymbol{\eta}},t+1}=\begin{pmatrix}\Delta{\mathbf{M}}_{t+1}\\ \Delta{\mathbf{M}}_{t+1}\end{pmatrix},\qquad R=\begin{pmatrix}UDV^{\top}&\mathcal{O}\\ \mathcal{O}&UV^{\top}\end{pmatrix}$.
Now, let us define the $(2N)\times(2N-1)$ matrices
$U_{\boldsymbol{\eta}}=\begin{pmatrix}U&{\boldsymbol{0}}&0\\ 0&{\mathbf{u}}&U\\ \end{pmatrix},\qquad\mbox{and}\qquad V_{\boldsymbol{\eta}}^{\top}=\begin{pmatrix}V^{\top}&0\\ {\boldsymbol{0}}^{\top}&{\mathbf{v}}^{\top}\\ 0&V^{\top}\\ \end{pmatrix}$,
so that $V_{\boldsymbol{\eta}}^{\top}U_{\boldsymbol{\eta}}=I$ and
$U_{\boldsymbol{\eta}}V_{\boldsymbol{\eta}}^{\top}=\begin{pmatrix}UV^{\top}&0\\ {\boldsymbol{0}}^{\top}&I\end{pmatrix}$.
Then, defining the $(2N-1)\times(2N-1)$ matrices
$S_{Q}=\begin{pmatrix}I\phi^{*}-D&{\boldsymbol{0}}&0\\ {\boldsymbol{0}}^{\top}&\phi^{*}&{\boldsymbol{0}}^{\top}\\ -I&{\boldsymbol{0}}&I\phi^{*}\end{pmatrix},\qquad\mbox{and}\qquad S_{R}=\begin{pmatrix}D&{\boldsymbol{0}}&0\\ {\boldsymbol{0}}^{\top}&0&{\boldsymbol{0}}^{\top}\\ 0&{\boldsymbol{0}}&I\end{pmatrix}$,
we have that $Q=U_{\boldsymbol{\eta}}S_{Q}V_{\boldsymbol{\eta}}^{\top}$ and $R=U_{\boldsymbol{\eta}}S_{R}V_{\boldsymbol{\eta}}^{\top}$. Then, from the above relations on $U_{\boldsymbol{\eta}}$ and $V_{\boldsymbol{\eta}}^{\top}$ we get $U_{\boldsymbol{\eta}}V_{\boldsymbol{\eta}}^{\top}{\boldsymbol{\eta}}_{t}={\boldsymbol{\eta}}_{t}$, and so multypling the dynamics of ${\boldsymbol{\eta}}_{t}$ by $U_{\boldsymbol{\eta}}V_{\boldsymbol{\eta}}^{\top}$ we finally get
${\boldsymbol{\eta}}_{t+1}=\frac{1}{t+1}U_{\boldsymbol{\eta}}\left(I-\frac{1}{t+1}S_{Q}\right)V_{\boldsymbol{\eta}}^{\top}{\boldsymbol{\eta}}_{t}+\frac{1}{t+1}R\Delta{\mathbf{M}}_{{\boldsymbol{\eta}},t+1}+{\mathbf{R}}_{{\boldsymbol{\eta}},t+1}.$
Step 2 (Explicit expression for ${\boldsymbol{\eta}}_{t}=(\widehat{{\mathbf{A}}}_{t},\widehat{{\mathbf{B}}}_{t})$).
Denoting the eigenvalues of $\Phi^{\top}$ different from $\phi^{*}$ by $\phi_{j}$ for $j=2,\dots,N$, let us set $\alpha_{j}=\phi^{*}-\phi_{j}$. Recall that $\mathcal{R}e(\alpha_{j})>0$ since $\phi^{*}$ is the leading eigenvalue and so $\mathcal{R}e(\phi_{j})<\phi^{*}$. Then, setting an initial time $m_{0}$ such that ${\mathcal{R}e}(\alpha_{j})/(m_{0}+1)<1$ for each $j$, we can write
${\boldsymbol{\eta}}_{t+1}$ $=$ $\ C_{m_{0},t}{\boldsymbol{\eta}}_{m_{0}}+\sum_{k=m_{0}}^{t}\frac{1}{k+1}C_{k+1,t}R\Delta{\mathbf{M}}_{{\boldsymbol{\eta}},k+1}+\sum_{k=m_{0}}^{t}C_{k+1,t}{\mathbf{R}}_{{\boldsymbol{\eta}},k+1}$
$=$ $\ C_{m_{0},t}{\boldsymbol{\eta}}_{m_{0}}+\sum_{k=m_{0}}^{t}{\mathbf{T}}_{t,k+1}+\boldsymbol{{\rho}}_{t+1}\,$,
where
(A.5) $\begin{split}C_{k+1,t}&=U_{\boldsymbol{\eta}}A_{k+1,t}V_{\boldsymbol{\eta}}^{\top},\qquad\mbox{and}\\ A_{k+1,t}&=\begin{cases}\prod_{m=k+1}^{t}\left(I-\frac{1}{m+1}S_{Q}\right)=\begin{pmatrix}A^{11}_{k+1,t}&{\boldsymbol{0}}&0\\ 0&a^{22}_{k+1,t}&0\\ A^{31}_{k+1,t}&{\boldsymbol{0}}&A^{33}_{k+1,t}\\ \end{pmatrix}\quad&k=m_{0}-1,\dots,t-1,\\ I\quad&k=t\,.\end{cases}\end{split}$
Notice that the blocks $A^{11}_{k+1,t}$, $A^{31}_{k+1,t}$ and $A^{33}_{k+1,t}$ are all diagonal $(N-1)\times(N-1)$ matrices. Moreover, setting for any $x\in\mathbb{C}$ with ${\mathcal{R}e}(x)/(m_{0}+1)<1$, $p_{m_{0}-1}(x)=1$ and $p_{k}(x)=\prod_{m=m_{0}}^{k}(1-\tfrac{x}{m+1})$ for $k\geq m_{0}$ and $F_{k+1,t}(x)=\frac{p_{t}(x)}{p_{k}(x)}$ for $m_{0}-1\leq k\leq t-1$, we get (see the supplementary material of Aletti et al. (2019) for similar technical calculations)
$j$ $=$ $\ F_{k+1,t}(\alpha_{j})$
$[A^{33}_{k+1,t}]_{jj}$ $=$ $\ a^{22}_{k+1,t}\ =\ F_{k+1,t}(\phi^{*})$
$[A^{31}_{k+1,t}]_{jj}$ $=$ $\ \left\{\begin{aligned} &-\frac{F_{k+1,t}(\alpha_{j})-F_{k+1,t}(\phi^{*})}{\alpha_{j}-\phi^{*}}&\mbox{for }\alpha_{j}\neq\phi^{*}\mbox{ ,i.e. }\phi_{j}\neq 0\\ &\frac{1-\phi^{*}}{\phi^{*}}F_{k+1,t}(\phi^{*})\ln\left(\frac{t}{k}\right)+O\left(\frac{1}{t}\right)&\mbox{for }\alpha_{j}=\phi^{*}\mbox{ ,i.e. }\phi_{j}=0.\end{aligned}\right$.
Step 3 (Study of the terms $C_{m_{0},t}{\boldsymbol{\eta}}_{m_{0}}$ and $\boldsymbol{{\rho}}_{t+1}$).
We will prove that $t^{\frac{\phi^{*}}{2}}|C_{m_{0},t}{\boldsymbol{\eta}}_{m_{0}}|\stackrel{{\scriptstyle a.s.}}{{\to}}0$ and $t^{\frac{\phi^{*}}{2}}|\boldsymbol{{\rho}}_{t+1}|\stackrel{{\scriptstyle a.s.}}{{\to}}0$. To this end, first notice that $O(|C_{k+1,t}|)=O(|A_{k+1,t}|)$ and, setting $a_{2}^{*}=\phi^{*}-{\mathcal{R}e}(\phi^{*}_{2})$, we have that
$|A_{k+1,t}|$ $=$ $\ O\left(\frac{|p_{t}(\alpha_{2}^{*})|}{|p_{k}(\alpha_{2}^{*})|}\right)+O\left(\frac{|p_{t}(\phi^{*})|}{|p_{k}(\phi^{*})|}\right)+O\left(\frac{|p_{t}(\phi^{*})|}{|p_{k}(\phi^{*})|}\ln\left(\frac{t}{k}\right)\right)+O\left(\frac{1}{t}\right)$
$=$ $\ O\left(\left(\frac{k}{t}\right)^{a_{2}^{*}}\right)+O\left(\left(\frac{k}{t}\right)^{\phi^{*}}\right)+O\left(\left(\frac{k}{t}\right)^{\phi^{*}}\ln\left(\frac{t}{k}\right)\right)+O\left(\frac{1}{t}\right)$
$=$ $\ O\left(\left(\frac{k}{t}\right)^{a_{2}^{*}}\right)+O\left(\left(\frac{k}{t}\right)^{\phi^{*}}\ln(t)\right)$
$=$ $\ O\left(\left(\frac{k}{t}\right)^{(\phi^{*}-{\mathcal{R}e}(\phi^{*}_{2})}\right)+O\left(\left(\frac{k}{t}\right)^{\phi^{*}}\ln(t)\right)\quad\mbox{for }k=m_{0}-1,\dots,t-1,$
and simply $|A_{t+1,t}|=O(1)$ for $k=t$. Therefore, since assumption $(v)$, we have
$t^{\frac{\phi^{*}}{2}}|C_{m_{0},t}{\boldsymbol{\eta}}_{m_{0}}|=|{\boldsymbol{\eta}}_{m_{0}}|\,O\left(\left(\frac{1}{t}\right)^{\phi^{*}/2-{\mathcal{R}e}(\phi_{2}^{*})}+\left(\frac{1}{t}\right)^{\phi^{*}/2}\ln(t)\right)\stackrel{{\scriptstyle a.s.}}{{\longrightarrow}}0\,.$
Moreover, recalling that ${\mathbf{R}}_{{\boldsymbol{\eta}},t+1}=O(t^{-(1+\beta)}){\boldsymbol{1}}$ for some $\beta>\frac{\phi^{*}}{2}$, we have
$t^{\frac{\phi^{*}}{2}}|\boldsymbol{{\rho}}_{t+1}|$ $=$ $\ t^{\frac{\phi^{*}}{2}}\left|\sum_{k=m_{0}}^{t}C_{k+1,t}{\mathbf{R}}_{{\boldsymbol{\eta}},k+1}\right|$
$=$ $\ t^{\frac{\phi^{*}}{2}}O\left(\sum_{k=m_{0}}^{t-1}\left(\frac{k}{t}\right)^{a_{2}^{*}}\frac{1}{k^{{1+\beta}}}+\sum_{k=m_{0}}^{t-1}\left(\frac{k}{t}\right)^{\phi^{*}}\ln(t)\frac{1}{k^{{1+\beta}}}\right)+t^{\phi^{*}/2}O(1/t^{1+\beta})$
$=$ $\ t^{\frac{\phi^{*}}{2}}O\left(\frac{1}{t^{a_{2}^{*}}}\sum_{k=m_{0}}^{t-1}k^{a_{2}^{*}-1-\beta}+\frac{\ln(t)}{t^{\phi^{*}}}\sum_{k=m_{0}}^{t-1}k^{\phi^{*}-1-\beta}\right)+O(1/t^{1+\beta-\phi^{*}/2}),$
which tends to zero because $\beta>\frac{\phi^{*}}{2}$ and $a_{2}^{*}>0$.
Step 4 (Study of the “delta-martingale” term).
We want to apply Theorem D.3. Therefore we have to check conditions (1) and (2) in Theorem D.3.
Since the relation $V_{\boldsymbol{\eta}}^{\top}U_{\boldsymbol{\eta}}=I$ implies $V_{\boldsymbol{\eta}}^{\top}R=S_{R}V_{\boldsymbol{\eta}}^{\top}$, we have that
$\sum_{k=m_{0}}^{t}(t^{\frac{\phi^{*}}{2}}{\mathbf{T}}_{t,k+1})(t^{\frac{\phi^{*}}{2}}{\mathbf{T}}_{t,k+1}^{\top})$ $=$ $\ t^{\phi^{*}}\sum_{k=m_{0}}^{t}\frac{1}{(k+1)^{2}}C_{k+1,t}R\Delta{\mathbf{M}}_{{\boldsymbol{\eta}},k+1}\Delta{\mathbf{M}}_{{\boldsymbol{\eta}},k+1}^{\top}R^{\top}C_{k+1,t}^{\top}$
$=$ $\ U_{\boldsymbol{\eta}}\left(t^{\phi^{*}}\sum_{k=m_{0}}^{t}\frac{1}{(k+1)^{2}}A_{k+1,t}V_{\boldsymbol{\eta}}^{\top}R\Delta{\mathbf{M}}_{{\boldsymbol{\eta}},k+1}\Delta{\mathbf{M}}_{{\boldsymbol{\eta}},k+1}^{\top}R^{\top}V_{\boldsymbol{\eta}}A^{\top}_{k+1,t}\right)U_{\boldsymbol{\eta}}^{\top}$
$=$ $\ U_{\boldsymbol{\eta}}\left(t^{\phi^{*}}\sum_{k=m_{0}}^{t}\frac{1}{(k+1)^{2}}A_{k+1,t}S_{R}V_{\boldsymbol{\eta}}^{\top}\Delta{\mathbf{M}}_{{\boldsymbol{\eta}},k+1}\Delta{\mathbf{M}}_{{\boldsymbol{\eta}},k+1}^{\top}V_{\boldsymbol{\eta}}S_{R}A^{\top}_{k+1,t}\right)U_{\boldsymbol{\eta}}^{\top}.$
Therefore, omitting the last term of the sum because $|A_{t+1,t}|=O(1)$ and , by assumption $(iii)$, we have $(t^{\phi^{*}}/(t+1)^{2})|\Delta{\mathbf{M}}_{t+1}|=O(t^{\phi^{*}-2+1-\eta})=O(t^{-(1-\phi^{*}+\eta)})\to 0$, it is enough to study the convergence of
(A.6) $t^{\phi^{*}}\sum_{k=m_{0}}^{t-1}\frac{1}{(k+1)^{2}}A_{k+1,t}S_{R}V_{\boldsymbol{\eta}}^{\top}\Delta{\mathbf{M}}_{{\boldsymbol{\eta}},k+1}\Delta{\mathbf{M}}_{{\boldsymbol{\eta}},k+1}^{\top}V_{\boldsymbol{\eta}}S_{R}A^{\top}_{k+1,t}.$
To this purpose, setting $B_{{\boldsymbol{\eta}},k+1}=V_{\boldsymbol{\eta}}^{\top}\Delta{\mathbf{M}}_{{\boldsymbol{\eta}},k+1}\Delta{\mathbf{M}}_{{\boldsymbol{\eta}},k+1}^{\top}V_{\boldsymbol{\eta}}$, $B_{k+1}=V^{\top}\Delta{\mathbf{M}}_{k+1}\Delta{\mathbf{M}}_{k+1}^{\top}V$, ${\mathbf{b}}_{k+1}=V^{\top}\Delta{\mathbf{M}}_{k+1}\Delta{\mathbf{M}}_{k+1}^{\top}{\mathbf{v}}$ and $b_{k}={\mathbf{v}}^{\top}\Delta{\mathbf{M}}_{k+1}\Delta{\mathbf{M}}_{k+1}^{\top}{\mathbf{v}}$, we observe that
$B_{{\boldsymbol{\eta}},k+1}=\begin{pmatrix}B_{k+1}&{\mathbf{b}}_{k+1}&B_{k+1}\\ {\mathbf{b}}^{\top}_{k+1}&b_{k+1}&{\mathbf{b}}^{\top}_{k+1}\\ B_{k+1}&{\mathbf{b}}_{k+1}&B_{k+1}\end{pmatrix}$.
Since in $B_{{\boldsymbol{\eta}},k+1}$ the first and the third row and column of blocks are the same, the matrix $(A_{k+1,t}S_{R})$ can be replaced by a diagonal matrix, with the following diagonal blocks: $A_{k+1,t}^{1}=A_{k+1,t}^{11}D$, $A_{k+1,t}^{3}=A_{k+1,t}^{31}D+A_{k+1,t}^{33}$ and $a_{k+1,t}^{2}=0$. Hence, the above expression (A.6) can be rewritten as
$t^{\phi^{*}}\sum_{k=m_{0}}^{t-1}\frac{1}{(k+1)^{2}}\begin{pmatrix}A_{k+1,t}^{1}B_{k+1}A_{k+1,t}^{1}&{\boldsymbol{0}}&A_{k+1,t}^{1}B_{k+1}A_{k+1,t}^{3}\\ {\boldsymbol{0}}^{\top}&0&{\boldsymbol{0}}^{\top}\\ A_{k+1,t}^{3}B_{k+1}A_{k+1,t}^{1}&{\boldsymbol{0}}&A_{k+1,t}^{3}B_{k+1}A_{k+1,t}^{3}\end{pmatrix}$.
The elements of $A_{k+1,t}^{1}$ and $A_{k+1,t}^{3}$ in the above matrix can be rewritten in terms of $F_{k+1,t}(\cdot)$ as
$j$ $=$ $\ (\phi^{*}-\alpha_{j})F_{k+1,t}(\alpha_{j})=\phi_{j}F_{k+1,t}(\alpha_{j})$
[…]
Morever, with some computations (see the supplementary material of Aletti et al. (2020) for similar technical calculations) for complex numbers $x$, $y$ such that ${\mathcal{R}e}(x+y)-\phi^{*}>0$ we have
(A.7) $t^{\phi^{*}}\sum_{k=m_{0}}^{t-1}\frac{1}{(k+1)^{1+\phi^{*}}}F_{k+1,t}(x)F_{k+1,t}(y)\rightarrow\frac{1}{x+y-\phi^{*}}.$
Hence, combining (see (Crimaldi et al., 2016, Lemma 4.1)) this limit with the limit relation
$t^{-(1-\phi^{*})}E[B_{t+1}|\mathcal{F}_{t}]=t^{-(1-\phi^{*})}V^{\top}E[\Delta{\mathbf{M}}_{t+1}\Delta{\mathbf{M}}_{t+1}^{\top}|\mathcal{F}_{t}]V\rightarrow V^{\top}\Sigma_{\infty}V,$
we get
(A.8) $t^{\phi^{*}}\sum_{k=m_{0}}^{t-1}\frac{1}{(k+1)^{1+\phi^{*}}}\left((k+1)^{-(1-\phi^{*})}B_{k+1}\right)F_{k+1,t}(x)F_{k+1,t}(y)\rightarrow\frac{1}{x+y-\phi^{*}}V^{\top}\Sigma_{\infty}V\,.$
As a consequence, recalling that $\alpha_{j}=\phi^{*}-\phi_{j}$ and, by assumption $(v)$, $\phi^{*}-2{\mathcal{R}e}(\phi^{*}_{2})>0$, the almost sure convergences of all the elements of the previous matrix can be obtained by applying (A.8) as follows:
$t^{\phi^{*}}\sum_{k=m_{0}}^{t-1}\frac{1}{(k+1)^{2}}[A_{k+1,t}^{1}B_{k+1}A_{k+1,t}^{1}]_{h,j}\rightarrow[{\mathcal{S}}_{\infty}^{11}]_{h,j}=\frac{\phi_{j}\phi_{h}}{\phi^{*}-\phi_{j}-\phi_{h}}({\mathbf{v}}_{j}^{\top}\Sigma_{\infty}{\mathbf{v}}_{h}),$
$t^{\phi^{*}}\sum_{k=m_{0}}^{t-1}\frac{1}{(k+1)^{2}}[A_{k+1,t}^{3}B_{k+1}A_{k+1,t}^{3}]_{h,j}\rightarrow[{\mathcal{S}}_{\infty}^{33}]_{h,j}=\frac{1}{\phi^{*}-\phi_{j}-\phi_{h}}({\mathbf{v}}_{j}^{\top}\Sigma_{\infty}{\mathbf{v}}_{h}),$
$t^{\phi^{*}}\sum_{k=m_{0}}^{t-1}\frac{1}{(k+1)^{2}}[A_{k+1,t}^{1}B_{k+1}A_{k+1,t}^{3}]_{h,j}\rightarrow[{\mathcal{S}}_{\infty}^{13}]_{h,j}=\frac{\phi_{j}}{\phi^{*}-\phi_{j}-\phi_{h}}({\mathbf{v}}_{j}^{\top}\Sigma_{\infty}{\mathbf{v}}_{h}).$
Therefore condition (1) of Theorem D.3 is satisfied with the matrix $\mathcal{M}$$_{\infty}$ defined in (A.3).
Regarding condition (2) of Theorem D.3, we observe that, by assumption $(iii)$, we have
$\begin{split}|{\mathbf{T}}_{t,k+1}|&=O\left(\frac{1}{k+1}|A_{k+1,t}||\Delta{\mathbf{M}}_{{\boldsymbol{\eta}},k+1}|\right)=O\left(|A_{k+1,t}|(k+1)^{-\eta}\right)\\ &=O\left(k^{-\eta}\,\left(\frac{k}{t}\right)^{(\phi^{*}-{\mathcal{R}e}(\phi^{*}_{2})}\right)+O\left(k^{-\eta}\,\left(\frac{k}{t}\right)^{\phi^{*}}\ln(t)\right)\quad\mbox{for }k=m_{0},\dots,t-1\end{split}$
and $|{\mathbf{T}}_{t,t+1}|=O\left(\frac{1}{t+1}|A_{t+1,t}||\Delta{\mathbf{M}}_{{\boldsymbol{\eta}},t+1}|\right)=O(t^{-\eta})$ for $k=t$. Hence, for any $u$, we have
$\left(\sup_{m_{0}\leq k\leq t}\Big{|}t^{\phi^{*}/2}{{\mathbf{T}}}_{t,k+1}\Big{|}\right)^{2u}$ $\leq t^{\phi^{*}u}\sum_{k=m_{0}}^{t-1}|{{\mathbf{T}}}_{t,k+1}|^{2u}+t^{\phi^{*}u}|{{\mathbf{T}}}_{t,t+1}|^{2u}$
$=O\left(t^{-(\phi^{*}-2{\mathcal{R}e}(\phi^{*}_{2}))u}\sum_{k=m_{0}}^{t-1}k^{-(\eta-\phi^{*}+2{\mathcal{R}e}(\phi^{*}_{2}))u}\right)$
$+O\left(\frac{\ln(t)^{2u}}{t^{\phi^{*}u}}\sum_{k=m_{0}}^{t-1}k^{-(\eta-\phi^{*})2u}\right)+O(t^{-(2\eta-\phi^{*})u}),$
which tends to zero for $u>\max\{1/\eta;1/(2\eta-\phi^{*})\}$ since $\phi^{*}-2{\mathcal{R}e}(\phi^{*}_{2})>0$ and $\eta>\phi^{*}/2$. This result, in particular, implies condition $(2)$ of Theorem D.3.
By Theorem D.3, we can finally conclude that $t^{\phi^{*}/2}\sum_{k=m_{0}}^{t}{\mathbf{T}}_{t,k+1}$ converges stably to the Gaussian kernel ${\mathcal{N}}({\boldsymbol{0}},{\mathcal{M}}_{\infty})$. ∎
Remark A.1.
A key point in the proof of Theorem A.3 is the convergence of the left-hand term of (A.7) toward a bounded limit, which is ensured by assumption (v), i.e. ${\mathcal{R}e}(\phi_{2}^{*})<\phi^{*}/{2}$. On the other hand, when ${\mathcal{R}e}(\phi_{2}^{*})\geq\phi^{*}/2$, this quantity can diverge to $\infty$, implying the divergence of $t^{\phi^{*}/2}(\widehat{{\mathbf{A}}}_{t},\widehat{{\mathbf{B}}}_{t})$. In fact, since $\Phi$ is a real matrix, in its spectrum we have $\phi_{2}^{*}$ and its conjugate, so that $F_{k+1,t}(x)F_{k+1,t}(y)=|F_{k+1,t}(x)|^{2}$ when we take $x=\alpha_{2}^{*}=\phi^{*}-\phi_{2}^{*}$ and for $y$ its conjugate (or $y=x$ if $x$ is real). Now, we observe that $|F_{k+1,t}(\alpha_{2}^{*})|^{2}$ increases with ${\mathcal{R}e}(\phi_{2}^{*})$ and the corresponding limit in (A.7), that is $1/(2{\mathcal{R}e}(\alpha_{2}^{*})-\phi^{*})=1/(\phi^{*}-2{\mathcal{R}e}(\phi_{2}^{*}))$, goes to $\infty$ for ${\mathcal{R}e}(\phi_{2}^{*})\uparrow\phi^{*}/2$. As a consequence of these two facts, we can conclude that the limit of the left-hand term of (A.7) is necessarily $\infty$ whenever ${\mathcal{R}e}(\phi_{2}^{*})\geq\phi^{*}/2$.
Theorem A.4.
When $N=1$, under assumptions $(i)$, $(ii)$, $(iii)$ and $(iv)$, we have
$t^{\phi^{*}/2}\begin{pmatrix}A_{t}-\widetilde{A}_{\infty}\\ B_{t}-\widetilde{A}_{\infty}/\phi^{*}\end{pmatrix}\stackrel{{\scriptstyle stably}}{{\longrightarrow}}{\mathcal{N}}\left(0,\begin{pmatrix}\phi^{*}\Sigma_{\infty}&\Sigma_{\infty}\\ \Sigma_{\infty}&\Sigma_{\infty}/\phi^{*}\end{pmatrix}\right)\,$.
When $N\geq 2$, under assumptions $(i)$, $(ii)$, $(iii)$, $(iv)$ and $(v)$, we have
[…]
where $\mathcal{M}$$_{\infty}$ is given in (A.3).
Proof.
When $N=1$, we simply have $(A_{t},B_{t})=(\widetilde{A}_{t},\widetilde{A}_{t}/\phi^{*}+o(t^{-\phi^{*}/2}))$ and so the statement follows from the clt for $\widetilde{A}_{t}$. When $N\geq 2$, we have
$\begin{split}&t^{\phi^{*}/2}({\mathbf{A}}_{t}-\widetilde{A}_{\infty}{\mathbf{u}},{\mathbf{B}}_{t}-\widetilde{A}_{\infty}/\phi^{*}{\mathbf{u}})\\ &=t^{\phi^{*}/2}((\widetilde{A}_{t}-\widetilde{A}_{\infty}){\mathbf{u}}+\widehat{{\mathbf{A}}}_{t},(\widetilde{A}_{t}-\widetilde{A}_{\infty})/\phi^{*}{\mathbf{u}}+\widehat{{\mathbf{B}}}_{t})\\ &=t^{\phi^{*}/2}(\widetilde{A}_{t}-\widetilde{A}_{\infty})({\mathbf{u}},{\mathbf{u}}/\phi^{*})+t^{\phi^{*}/2}(\widehat{{\mathbf{A}}}_{t},\widehat{{\mathbf{B}}}_{t})\,.\end{split}$
The first term converges in the sense of the almost sure conditional convergence toward the Gaussian kernel with mean $\boldsymbol{0}$} and covariance matrix
${\mathbf{v}}^{\top}\Sigma_{\infty}{\mathbf{v}}\begin{pmatrix}\phi^{*}{\mathbf{u}}{\mathbf{u}}^{\top}&{\mathbf{u}}{\mathbf{u}}^{\top}\\ {\mathbf{u}}{\mathbf{u}}^{\top}&{\mathbf{u}}{\mathbf{u}}^{\top}/\phi^{*}\end{pmatrix}\,$.
The second term converges stably toward the Gaussian kernel ${\mathcal{N}}({\boldsymbol{0}},{\mathcal{M}}_{\infty})$. By Lemma D.4, we can combine the two convergences and obtain the desidered convergence. ∎
As an immediate corollary, we can get the central limit theorem for $\widetilde{B}_{t}={\mathbf{v}}^{\top}{\mathbf{B}}_{t}$ when $N\geq 2$.
Corollary A.5.
(CLT for $(\widetilde{B}_{t})$ when $N\geq 2$)
When $N\geq 2$, under assumptions $(i)$, $(ii)$, $(iii)$, $(iv)$ and $(v)$, we have
$t^{\phi^{*}/2}(\widetilde{B}_{t}-\widetilde{A}_{\infty}/\phi^{*}){\stackrel{{\scriptstyle stably}}{{\longrightarrow}}}\mathcal{N}\left(0\ ,{\mathbf{v}}^{\top}{\Sigma}_{\infty}{\mathbf{v}}/\phi^{*}\right)\,.$
Proof.
Setting $\widetilde{B}_{t}={\mathbf{v}}^{\top}{\mathbf{B}}_{t}$ and using the relations ${\mathbf{v}}^{\top}{\mathbf{u}}=1$ and ${\mathbf{v}}^{\top}U={\boldsymbol{0}}$ in Theorem A.4, we find the desidered convergence. ∎
We can also derive the following result that consists in a CLT for a normalized version of $({\mathbf{A}}_{t},{\mathbf{B}}_{t})$.
Corollary A.6.
(CLT for $({\mathbf{A}}_{t}/({\mathbf{x}}^{\top}{\mathbf{A}}_{t}),{\mathbf{B}}_{t}/({\mathbf{x}}^{\top}{\mathbf{B}}_{t}))$ when $N\geq 2$)
Under assumptions $(i)$, $(ii)$, $(iii)$, $(iv)$ and $(v)$, if $\widetilde{A}_{\infty}\neq 0$ a.s., then, for any given vector with ${\mathbf{x}}^{\top}{\mathbf{u}}\neq 0$, we have
$t^{\phi^{*}/2}\begin{pmatrix}{\mathbf{A}}_{t}/({\mathbf{x}}^{\top}{\mathbf{A}}_{t})-{\mathbf{u}}/({\mathbf{x}}^{\top}{\mathbf{u}})\\ {\mathbf{B}}_{t}/({\mathbf{x}}^{\top}{\mathbf{B}}_{t})-{\mathbf{u}}/({\mathbf{x}}^{\top}{\mathbf{u}})\end{pmatrix}\stackrel{{\scriptstyle stably}}{{\longrightarrow}}{\mathcal{N}}\left({\boldsymbol{0}},\frac{1}{({\mathbf{x}}^{\top}{\mathbf{u}})^{2}\widetilde{A}_{\infty}^{2}}\begin{pmatrix}Q_{\mathbf{x}}\mathcal{M}_{\infty}^{11}Q_{\mathbf{x}}^{\top}&\phi^{*}Q_{\mathbf{x}}\mathcal{M}_{\infty}^{13}Q_{\mathbf{x}}^{\top}\\ \phi^{*}Q_{\mathbf{x}}\mathcal{M}_{\infty}^{13}Q_{\mathbf{x}}^{\top}&(\phi^{*})^{2}Q_{\mathbf{x}}\mathcal{M}_{\infty}^{33}Q_{\mathbf{x}}^{\top}\end{pmatrix}\right)$,
where $Q_{\mathbf{x}}=I-{\mathbf{u}}{\mathbf{x}}^{\top}/({\mathbf{x}}^{\top}{\mathbf{u}})$ and the matrices $\mathcal{M}_{\infty}^{11},\,\mathcal{M}_{\infty}^{13},\,\mathcal{M}_{\infty}^{33}$ are defined in (A.3). In particular, for ${\mathbf{x}}={\mathbf{v}}$ we have
$t^{\phi^{*}/2}\begin{pmatrix}{\mathbf{A}}_{t}/\widetilde{A}_{t}-{\mathbf{u}}\\ {\mathbf{B}}_{t}/\widetilde{B}_{t}-{\mathbf{u}}\end{pmatrix}\stackrel{{\scriptstyle stably}}{{\longrightarrow}}{\mathcal{N}}\left({\boldsymbol{0}},\frac{1}{\widetilde{A}_{\infty}^{2}}\begin{pmatrix}\mathcal{M}_{\infty}^{11}&\phi^{*}\mathcal{M}_{\infty}^{13}\\ \phi^{*}\mathcal{M}_{\infty}^{13}&(\phi^{*})^{2}\mathcal{M}_{\infty}^{33}\end{pmatrix}\right)\,$.
Remark A.2.
Note that, by (A.2) and (A.4), we have $\widehat{{\mathbf{A}}}_{t}={\mathbf{A}}_{t}-\widetilde{A}_{t}{\mathbf{u}}=UV^{\top}{\mathbf{A}}_{t}$ and, similarly, ${\mathbf{B}}_{t}-\widetilde{B}_{t}{\mathbf{u}}=UV^{\top}{\mathbf{B}}_{t}$. Therefore, the processes in the above result in the case ${\mathbf{x}}={\mathbf{v}}$ coincide, up to multiplicative (random) factors, with the processes $UV^{\top}{\mathbf{A}}_{t}$ and $UV^{\top}{\mathbf{B}}_{t}$ (note also that we have $Q_{{\mathbf{v}}}=I-{\mathbf{u}}{\mathbf{v}}^{\top}/({\mathbf{v}}^{\top}{\mathbf{u}})=UV^{\top}$ by (A.4)). These processes will play a fundamental role in the following section dedicated to the statistical tools.
Proof.
The result follows from the CLT of Theorem A.4 once we note that
$t^{\phi^{*}/2}$ $\begin{pmatrix}{\mathbf{A}}_{t}/({\mathbf{x}}^{\top}{\mathbf{A}}_{t})-{\mathbf{u}}/({\mathbf{x}}^{\top}{\mathbf{u}})\\ {\mathbf{B}}_{t}/({\mathbf{x}}^{\top}{\mathbf{B}}_{t})-{\mathbf{u}}/({\mathbf{x}}^{\top}{\mathbf{u}})\end{pmatrix}$
$=t^{\phi^{*}/2}\begin{pmatrix}({\mathbf{x}}^{\top}{\mathbf{A}}_{t})^{-1}I&0\\ 0&({\mathbf{x}}^{\top}{\mathbf{B}}_{t})^{-1}I\\ \end{pmatrix}\begin{pmatrix}{\mathbf{A}}_{t}-{\mathbf{u}}{\mathbf{x}}^{\top}{\mathbf{A}}_{t}/({\mathbf{x}}^{\top}{\mathbf{u}})\\ {\mathbf{B}}_{t}-{\mathbf{u}}{\mathbf{x}}^{\top}{\mathbf{B}}_{t}/({\mathbf{x}}^{\top}{\mathbf{u}})\end{pmatrix}$
$=t^{\phi^{*}/2}\begin{pmatrix}({\mathbf{x}}^{\top}{\mathbf{A}}_{t})^{-1}Q_{\mathbf{x}}&0\\ 0&({\mathbf{x}}^{\top}{\mathbf{B}}_{t})^{-1}Q_{\mathbf{x}}\\ \end{pmatrix}\begin{pmatrix}{\mathbf{A}}_{t}\\ {\mathbf{B}}_{t}\end{pmatrix}$
$=t^{\phi^{*}/2}\begin{pmatrix}({\mathbf{x}}^{\top}{\mathbf{A}}_{t})^{-1}Q_{\mathbf{x}}&0\\ 0&({\mathbf{x}}^{\top}{\mathbf{B}}_{t})^{-1}Q_{\mathbf{x}}\\ \end{pmatrix}\begin{pmatrix}{\mathbf{A}}_{t}-\widetilde{A}_{\infty}{\mathbf{u}}\\ {\mathbf{B}}_{t}-\widetilde{A}_{\infty}/\phi^{*}{\mathbf{u}}\end{pmatrix}$.
Indeed, we have $Q_{\mathbf{x}}{\mathbf{u}}{\mathbf{u}}^{\top}=0$ and so
[…]
The result for ${\mathbf{x}}={\mathbf{v}}$ simply follows by the fact that, since (A.4), we have the equalities $Q_{\mathbf{x}}=Q_{{\mathbf{v}}}=(I-{\mathbf{u}}{\mathbf{v}}^{\top}/({\mathbf{v}}^{\top}{\mathbf{u}}))=UV^{\top}$ and so $UV^{\top}{\mathcal{M}}_{\infty}^{ij}VU^{\top}={\mathcal{M}}_{\infty}^{ij}$. ∎
Appendix B Statistical tools
In this section we present some statistical tools, based on Theorem A.4, to provide an interval estimation of the random limit $\widetilde{A}_{\infty}$ and an hypothesis test on the structure of the matrix $\Phi$. To this end, we need to assume that the asymptotic conditional covariance matrix $\Sigma_{\infty}$ of the martingale vector $\Delta{\mathbf{M}}_{t}$ satisfies the following assumption:
Assumption B.1.
Assume the matrix $\Sigma_{\infty}$ can be factorized as
$\Sigma_{\infty}=g(\widetilde{A}_{\infty})\Sigma_{det}\,$,
with $g$ being a continuous function with $g(\widetilde{A}_{\infty})>0$ a.s. and $\Sigma_{det}$ a deterministic positive definite symmetric matrix determined by the eigen-structure of $\Phi$.
We will provide statistical tools based on the process $({\mathbf{A}}_{t})$ or on the process $({\mathbf{B}}_{t})$ so that the suitable tool can be selected in the practical applications when only one of the two processes is observable.
B.1. Confidence intervals for the random limit $\widetilde{A}_{\infty}$ based on $({\mathbf{A}}_{t})$
First of all, we can observe that by means of the central limit theorem for the stochastic process $\widetilde{A}_{t}={{\mathbf{v}}}^{\top}\,{{\mathbf{A}}}_{t}$ described in Theorem A.1, it is possible to construct an asymptotic confidence interval for its limit $\widetilde{A}_{\infty}$. Specifically, an asymptotic confidence interval for $\widetilde{A}_{\infty}$ with approximate level $(1-\alpha)$ is the following:
$CI_{1-\alpha}(\widetilde{A}_{\infty})\ :=\ \left(\ \widetilde{A}_{t}-\frac{z_{\alpha}}{t^{\phi^{*}/2}}\sqrt{g(\widetilde{A}_{t})\phi^{*}{\mathbf{v}}^{\top}\Sigma_{det}{\mathbf{v}}}\ ;\ \widetilde{A}_{t}+\frac{z_{\alpha}}{t^{\phi^{*}/2}}\sqrt{g(\widetilde{A}_{t})\phi^{*}{\mathbf{v}}^{\top}\Sigma_{det}{\mathbf{v}}}\ \right),$
[…]
B.2. Confidence intervals for the random limit $\widetilde{A}_{\infty}$ based on $({\mathbf{B}}_{t})$
We here construct an asymptotic confidence interval for $\widetilde{A}_{\infty}$ based on the process $({\mathbf{B}}_{t})_{t}$. First notice that using Theorem A.4 we can obtain for $N=1$
$CI_{1-\alpha}(\widetilde{A}_{\infty})\ =\ \phi^{*}B_{t}\ \pm\ \frac{z_{\alpha}}{t^{\phi^{*}/2}}\sqrt{g(\phi^{*}B_{t})\phi^{*}\Sigma_{det}},$
[…]
When $N\geq 2$, it is required the knowledge of ${\mathcal{R}e}(\phi_{2}^{*})$. Indeed, when ${\mathcal{R}e}(\phi_{2}^{*})/\phi^{*}<1/2$, setting $\widetilde{B}_{t}={{\mathbf{v}}}^{\top}{{\mathbf{B}}}_{t}$ and using Corollary A.5, we obtain
$CI_{1-\alpha}(\widetilde{A}_{\infty})\ =\ \phi^{*}\widetilde{B}_{t}\ \pm\ \frac{z_{\alpha}}{t^{\phi^{*}/2}}\!\sqrt{g(\phi^{*}\widetilde{B}_{t})\phi^{*}{\mathbf{v}}^{\top}\Sigma_{det}{\mathbf{v}}},$
where $z_{\alpha}$ is such that ${\mathcal{N}}(0,1)(z_{\alpha},+\infty)=\alpha/2$.
B.3. Hypothesis tests on the matrix $\Phi$ based on $({\mathbf{A}}_{t})$
We now focus on the case $N\geq 2$ and the inferential problem of testing the null hypothesis $H_{0}:\,\Phi=\Phi_{0}$, using the multi-dimensional stochastic process $({\mathbf{A}}_{t})_{t}$. Since the distribution of $\widetilde{A}_{\infty}$ should be unknown in practice, we propose a test statistics whose limit does not involve $\widetilde{A}_{\infty}$. First, we need to introduce some notation. Let us define the matrix ${\mathcal{M}}_{det}^{11}$ as the matrix $\mathcal{M}_{\infty}^{11}$ defined in Theorem A.3 with $\Sigma_{\infty}$ replaced by $\Sigma_{det}$, i.e. $\mathcal{M}_{\infty}^{11}=g(\widetilde{A}_{\infty}){\mathcal{M}}_{det}^{11}$. Then, notice that the matrix ${\mathcal{M}}_{det}^{11}$ has rank (N1)(N-1) and it admits a spectral decomposition as follows: ${\mathcal{M}}_{det}^{11}=O^{11}{\mathcal{D}}^{11}(O^{11})^{\top}$, where ${\mathcal{D}}^{11}$ is the diagonal matrix of dimension (N1)(N-1) containing the eigenvalues of ${\mathcal{M}}_{det}^{11}$ different from 0 and $O^{11}$ is the $N\times(N-1)$ matrix whose columns form an orthonormal basis of $Im(\mathcal{M}_{det}^{11})$ of corresponding eigenvectors. Then, we set $L^{11}=({\mathcal{D}}^{11})^{-1/2}$ (diagonal matrix of dimension (N1)(N-1)) and take the product $L^{11}(O^{11})^{\top}$ (matrix of dimension $(N-1)\times N$). Fixed the matrix assumed under $H_{0}$, i.e. $\Phi_{0}$, we can compute for it the eigenvectors $\mathbf{u}$}} and $\mathbf{v}$}}, the eigenvectors contained in the matrices $U$ and $V$ and the matrices $L^{11}$ and $O^{11}$ defined above. Hence, we can obtain under $H_{0}$ the real process $\widetilde{A}_{t}={{\mathbf{v}}}^{\top}\,{{\mathbf{A}}}_{t}$ and the multi-dimensional process $\widehat{{{\mathbf{A}}}}_{t}={\mathbf{A}}_{t}-\widetilde{A}_{t}{\mathbf{u}}=U\,V^{\top}\,{{\mathbf{A}}}_{t}$. Then, from Theorem A.3 we have that, under $H_{0}$ and the assumption ${\mathcal{R}e}(\phi_{2}^{*})/\phi^{*}<1/2$,
(B.1) ${{\mathbf{T}}}_{t}=t^{\phi^{*}/2}g(\widetilde{A}_{t})^{-1/2}L^{11}(O^{11})^{\top}\,U\,V^{\top}\,{{\mathbf{A}}}_{t}=t^{\phi^{*}/2}g(\widetilde{A}_{t})^{-1/2}L^{11}(O^{11})^{\top}\,\widehat{{{\mathbf{A}}}}_{t}\stackrel{{\scriptstyle d}}{{\longrightarrow}}\mathcal{N}\left({\boldsymbol{0}},I_{N-1}\right)$
(in the above formulas the matrices $L^{11}$ and $O^{11}$ are related to ${\mathcal{M}}_{det}^{11}$ computed under $H_{0}$ and so for $\Phi_{0}$). Hence, under $H_{0}$, the test statistics $\|{{\mathbf{T}}}_{t}\|^{2}$ is asymptotically chi-squared distributed with (N1)(N-1) degrees of freedom. This result lets us construct an asymptotic critical region for testing any $\Phi_{0}$.
Remark B.1.
[…]
$U_{0}V_{0}^{\top}{{\mathbf{A}}}_{t}=U_{0}V_{0}^{\top}({\mathbf{A}}_{t}-\widetilde{A}_{t}{\mathbf{u}}_{1})=U_{0}V_{0}^{\top}\widehat{{{\mathbf{A}}}}_{t}.$
In summary, whenever ${\mathbf{u}}_{0}={\mathbf{u}}_{1}$ we have under $H_{1}$ that ${{\mathbf{T}}}_{t}$ is still asymptotically normally distributed with covariance matrix $L^{11}(O^{11})^{\top}UV^{\top}{\mathcal{M}}_{det}^{11}VU^{\top}O^{11}(L^{11})^{\top}$, where ${\mathcal{M}}_{det}^{11}$ is referred to the eigen-structure of $\Phi_{1}$, while $L^{11}(O^{11})^{\top}UV^{\top}$ is referred to $\Phi_{0}$. Thus, the distance between the identity $I$ and $L^{11}(O^{11})^{\top}UV^{\top}{\mathcal{M}}_{det}^{11}VU^{\top}O^{11}(L^{11})^{\top}$ describes the relation between the asymptotic distribution of $\|{{\mathbf{T}}}_{t}\|^{2}$ under $H_{0}$ and the one under $H_{1}$, which determines the power of the test. For instance, note that $E[\|{{\mathbf{T}}}_{t}\|^{2}]=(N-1)$ under $H_{0}$, while $E[\|{{\mathbf{T}}}_{t}\|^{2}]$ is equal to the trace of $L^{11}(O^{11})^{\top}UV^{\top}{\mathcal{M}}_{det}^{11}VU^{\top}O^{11}(L^{11})^{\top}$ under $H_{1}$.
B.4. Hypothesis tests on the matrix $\Phi$ based on $({\mathbf{B}}_{t})$
We here want to build for the case $N\geq 2$ a test statistic for $H_{0}:\Phi=\Phi_{0}$, that is based only on the process $({\mathbf{B}}_{t})$. To this end, analogously as in Subsection B.3, let us define the matrix ${\mathcal{M}}_{det}^{33}$ such that $\mathcal{M}_{\infty}^{33}=g(\widetilde{A}_{\infty}){\mathcal{M}}_{det}^{33}$ and the corresponding matrices $L^{33}$ and $O^{33}$. Fixed the matrix $\Phi_{0}$ assumed under $H_{0}$, we can obtain the real process $\widetilde{B}_{t}={{\mathbf{v}}}^{\top}\,{{\mathbf{B}}}_{t}$ and the multi-dimensional process
${\mathbf{B}}^{\prime}_{t}={\mathbf{B}}_{t}-\widetilde{B}_{t}{\mathbf{u}}=U\,V^{\top}\,{{\mathbf{B}}}_{t}.$
By Corollary A.6 and Remark A.2, we get that, under $H_{0}$ and the assumption ${\mathcal{R}e}(\phi_{2}^{*})/\phi^{*}<1/2$, we have
(B.2) ${{\mathbf{T}}}_{t}=t^{\phi^{*}/2}g(\phi^{*}\widetilde{B}_{t})^{-1/2}L^{33}(O^{33})^{\top}{\mathbf{B}}^{\prime}_{t}=t^{\phi^{*}/2}g(\phi^{*}\widetilde{B}_{t})^{-1/2}L^{33}(O^{33})^{\top}\,U\,V^{\top}\,{{\mathbf{B}}}_{t}\stackrel{{\scriptstyle d}}{{\longrightarrow}}\mathcal{N}\left({\boldsymbol{0}},I_{N-1}\right)$
(in the above formulas the matrices $L^{33}$ and $O^{33}$ are related to ${\mathcal{M}}_{det}^{33}$ computed under $H_{0}$ and so for $\Phi_{0}$). The considerations on the power made for the test based on ${\mathbf{A}}_{t}$ in Remark B.1 remain valid as well for the test based on ${\mathbf{B}}_{t}$.
Appendix C Statistical tools in some examples
We here specify the previous general statistical tools in two particular cases: firstly, we consider the case $N=2$ and, then, we take into account the case when $\Phi$ is of the “mean-field” type and $N\geq 2$.
C.1. Case $N=2$
A matrix $\Phi$ with dimension $2\times 2$ and satisfying condition $(i)$ can be reparametrized imposing that the number $\phi\in(0,1]$ is an eigenvalue with left eigenvector $\mathbf{u}$} such that $r=u_{1}/u_{2}$, with $r>0$, that is
$r\phi_{11}+\phi_{21}=r\phi\quad\mbox{and}\quad r\phi_{12}+\phi_{22}=\phi\,.$
This reparametrization has been inspired by the fact that we typically are able to estimate with high accuracy the leading eigenvalue $\phi^{*}\in(0,1]$ and the ratio $r=u_{1}/u_{2}>0$ of the two componets of the leading left eigenvector $\mathbf{u}$}. Hence, we may think that $\phi=\phi^{*}$ and $r$ are known in practice and the only parameters that are object of the test are $\phi_{21}$ and $\phi_{12}$. We can also calculate the second eigenvalue from the trace of $\Phi$:
$\phi_{2}^{*}=\phi_{1,1}+\phi_{2,2}-\phi=\phi-\frac{\phi_{2,1}+r^{2}\phi_{1,2}}{r},$
so that the condition ${\mathcal{R}e}(\phi_{2}^{*})/\phi^{*}<1/2$ required in the previous results corresponds to the condition
$2\frac{\phi_{2,1}+r^{2}\phi_{1,2}}{r}>\phi.$
Moreover, since according to our notation we have ${\mathbf{v}}^{\top}{\boldsymbol{1}}=1$ and ${\mathbf{v}}^{\top}{\mathbf{u}}=1$, we obtain the following relations:
$v_{1}=\frac{r\phi_{1,2}}{r\phi_{1,2}+\phi_{2,1}},\qquad v_{2}=1-v_{1},\qquad u_{1}=\frac{r}{1-(1-r)v_{1}},\qquad u_{2}=\frac{1}{1-(1-r)v_{1}},$
which leads to
$v_{1}=\frac{r\phi_{1,2}}{r\phi_{1,2}+\phi_{2,1}},\qquad v_{2}=\frac{\phi_{2,1}}{r\phi_{1,2}+\phi_{2,1}},\qquad u_{1}=\frac{r\phi_{2,1}+r^{2}\phi_{1,2}}{\phi_{2,1}+r^{2}\phi_{1,2}},\qquad u_{2}=\frac{\phi_{2,1}+r\phi_{1,2}}{\phi_{2,1}+r^{2}\phi_{1,2}}.$
If we adopt the parametrization $\iota=\phi_{1,2}+\phi_{2,1}$ and $\eta=\tfrac{\phi_{1,2}}{\phi_{1,2}+\phi_{2,1}}$ so that $\iota$ represents the intensity of the whole “interaction” among the two components $1$ and $2$ due to matrix $\Phi$ and $\eta$ (resp. $(1-\eta)$) gives the percentage of this intensity due to the influence of the first component on the second one (resp., of the second component on the first one). Then, we have
${\mathbf{u}}=\frac{(1-\eta)+r\eta}{(1-\eta)+r^{2}\eta}\begin{pmatrix}r\\ 1\end{pmatrix}\qquad\mbox{and}\qquad{\mathbf{v}}=\frac{1}{r\eta+(1-\eta)}\begin{pmatrix}r\eta\\ 1-\eta\end{pmatrix}$,
and so
${\mathbf{u}}{\mathbf{v}}^{\top}=\frac{1}{(1-\eta)+r^{2}\eta}\begin{pmatrix}r\\ 1\end{pmatrix}\begin{pmatrix}r\eta&1-\eta\end{pmatrix}\qquad\mbox{and}\qquad UV^{\top}=\frac{1}{(1-\eta)+r^{2}\eta}\begin{pmatrix}1-\eta\\ -r\eta\end{pmatrix}\begin{pmatrix}1&-r\end{pmatrix}$.
Finally, imposing $U^{\top}U=1$, we get
$U=\frac{1}{\sqrt{(1-\eta)^{2}+r^{2}\eta^{2}}}\begin{pmatrix}1-\eta\\ -r\eta\end{pmatrix}\qquad\mbox{and}\qquad V=\frac{\sqrt{(1-\eta)^{2}+r^{2}\eta^{2}}}{(1-\eta)+r^{2}\eta}\begin{pmatrix}1\\ -r\end{pmatrix}$.
We thus obtain
$\phi^{*}{\mathbf{v}}^{\top}\Sigma_{det}{\mathbf{v}}=\frac{\phi}{(r\eta+(1-\eta))^{2}}\begin{pmatrix}r\eta&1-\eta\end{pmatrix}\Sigma_{det}\begin{pmatrix}r\eta\\ 1-\eta\end{pmatrix}$
and, for $2\iota\frac{(1-\eta)+r^{2}\eta}{r}>\phi$,
${\mathcal{M}}_{det}^{11}=U{\mathcal{S}}_{det}^{11}U^{\top}=\frac{(\iota\frac{(1-\eta)+r^{2}\eta}{r}-\phi)^{2}}{2\iota\frac{(1-\eta)+r^{2}\eta}{r}-\phi}\frac{1}{((1-\eta)+r^{2}\eta)^{2}}\begin{pmatrix}1-\eta\\ -r\eta\end{pmatrix}\begin{pmatrix}1&-r\end{pmatrix}\Sigma_{det}\begin{pmatrix}1\\ -r\end{pmatrix}\begin{pmatrix}1-\eta&-r\eta\end{pmatrix}$,
${\mathcal{M}}_{det}^{33}=U{\mathcal{S}}_{det}^{33}U^{\top}=\frac{1}{2\iota\frac{(1-\eta)+r^{2}\eta}{r}-\phi}\frac{1}{((1-\eta)+r^{2}\eta)^{2}}\begin{pmatrix}1-\eta\\ -r\eta\end{pmatrix}\begin{pmatrix}1&-r\end{pmatrix}\Sigma_{det}\begin{pmatrix}1\\ -r\end{pmatrix}\begin{pmatrix}1-\eta&-r\eta\end{pmatrix}$.
(where $\mathcal{S}_{det}^{ii}$ are the matrices defined in Theorem A.3 with $\Sigma_{\infty}$ replaced by $\Sigma_{det}$). Therefore, recalling that
$\widetilde{A}_{t}={\mathbf{v}}^{\top}{\mathbf{A}}_{t}=\frac{r\eta A_{1,t}+(1-\eta)A_{2,t}}{r\eta+(1-\eta)}\qquad\mbox{and}\qquad\widetilde{B}_{t}={\mathbf{v}}^{\top}{\mathbf{B}}_{t}=\frac{r\eta B_{1,t}+(1-\eta)B_{2,t}}{r\eta+(1-\eta)},$
and defining the constant
$c_{r,\eta}=\tfrac{\begin{pmatrix}r\eta&(1-\eta)\end{pmatrix}\Sigma_{det}\begin{pmatrix}r\eta\\ (1-\eta)\end{pmatrix}}{(r\eta+(1-\eta))^{2}},$
the confidence interval based on ${\mathbf{A}}_{t}$ and ${\mathbf{B}}_{t}$ are respectively
$CI_{1-\alpha}(\widetilde{A}_{\infty})\ =\ \widetilde{A}_{t}\pm\ \frac{z_{\alpha}}{t^{\phi/2}}\!\sqrt{\phi g(\widetilde{A}_{t})c_{r,\eta}}\,,\qquad\mbox{and}\qquad CI_{1-\alpha}(\widetilde{A}_{\infty})\ =\ \phi\widetilde{B}_{t}\pm\ \frac{z_{\alpha}}{t^{\phi/2}}\!\sqrt{\phi g(\phi\widetilde{B}_{t})c_{r,\eta}}.$
Analogously, the test statistics based on ${\mathbf{A}}_{t}$ is
$T_{t}=t^{\phi/2}\frac{1}{(1-\eta)+r^{2}\eta}L^{11}(O^{11})^{\top}\begin{pmatrix}(1-\eta)\\ -r\eta\end{pmatrix}\frac{(A_{1,t}-rA_{2,t})}{g(\widetilde{A}_{t})^{1/2}},$
and the one based on ${\mathbf{B}}_{t}$
$T_{t}=t^{\phi/2}\frac{1}{(1-\eta)+r^{2}\eta}L^{33}(O^{33})^{\top}\begin{pmatrix}(1-\eta)\\ -r\eta\end{pmatrix}\,\frac{(B_{1,t}-rB_{2,t})}{g(\phi\widetilde{B}_{t})^{1/2}}.$
Remark C.1.
In the particular case $\Sigma_{det}=diag({\mathbf{u}})$ we have
${\mathcal{S}}_{det}^{33}=\frac{1}{\left(2\iota\frac{(1-\eta)+r^{2}\eta}{r}-\phi\right)}\frac{(r+r^{2})((1-\eta)+r\eta)((1-\eta)^{2}+r^{2}\eta^{2})}{((1-\eta)+r^{2}\eta)^{3}},$
which leads to $L^{33}(O^{33})^{\top}=({\mathcal{S}}_{det}^{33})^{-1/2}U^{\top}$ and so the test statistics based on ${\mathbf{B}}_{t}$ becomes
$T_{t}=t^{\phi/2}\sqrt{\left(2\iota\frac{(1-\eta)+r^{2}\eta}{r}-\phi\right)\frac{(1-\eta)+r^{2}\eta}{(r+r^{2})((1-\eta)+r\eta)}}\,\frac{(B_{1,t}-rB_{2,t})}{g(\phi\widetilde{B}_{t})^{1/2}}.$
Then, denoting
$\Delta=\frac{1}{2}-\frac{\phi^{*}_{2}}{\phi}=\frac{\iota}{\phi}\frac{(1-\eta)+r^{2}\eta}{r}-\frac{1}{2}\qquad\mbox{and}\qquad q_{r,\eta}=\frac{2((1-\eta)+r^{2}\eta)}{r(1+r)((1-\eta)+r\eta)},$
we have
(C.1) $\|T_{t}\|^{2}=t^{\phi}q_{r,\eta}\Delta\,\frac{(B_{1,t}-rB_{2,t})^{2}}{g(\phi\widetilde{B}_{t})/\phi}\stackrel{{\scriptstyle d}}{{\longrightarrow}}\chi^{2}(N-1).$
Moreover, the confidence interval based on ${\mathbf{B}}_{t}$ becomes
(C.2) $CI_{1-\alpha}(\widetilde{A}_{\infty})\ =\ \phi\widetilde{B}_{t}\pm\ \frac{z_{\alpha}}{t^{\phi/2}}\!\sqrt{\phi g(\phi\widetilde{B}_{t})c_{r,\eta}}\,,\qquad\mbox{with}\qquad c_{r,\eta}=\tfrac{(r^{3}\eta^{2}+(1-\eta)^{2})}{(r\eta+(1-\eta))(r^{2}\eta+(1-\eta))}.$
C.2. Mean-field case for $N\geq 2$
This first case we consider is the “Mean-field” case, which refers to the family of matrices $\Phi$ that can be expressed as follows: for any $1\leq h,j\leq N$
$\phi_{j,h}\ =\ \phi\left(\frac{\iota}{N}\ +\ \delta_{j,h}(1-\iota)\right)\qquad\mbox{with }\phi,\iota\in(0,1],$
where $\delta_{j,h}$ is equal to $1$ when $h=j$ and to 0 otherwise. Note that $\Phi$ is irreducible and doubly stochastic, we have ${{\mathbf{v}}}=N^{-1}{{\boldsymbol{1}}}$ and ${{\mathbf{u}}}={\boldsymbol{1}}$ and so the random variable $\widetilde{A}_{t}$ coincides with the average of the processes $A_{t,h}$, i.e. $N^{-1}{\boldsymbol{1}}^{\top}{{\mathbf{A}}}_{t}$ and $\widehat{{{\mathbf{A}}}}_{t}=\left(I-N^{-1}{\boldsymbol{1}}{\boldsymbol{1}}^{\top}\right){{\mathbf{A}}}_{t}$. Furthermore, in this case, the leading eigenvalue $\phi^{*}$ coincides with $\phi$, while all the eigenvalues of $\Phi$ different from $\phi^{*}$ are equal to $\phi(1-\iota)$ and, consequently, the condition ${\mathcal{R}e}(\phi_{2}^{*})/\phi^{*}<1/2$ required in the previous results corresponds to the condition $2\iota>1$. Finally, since $\Phi$ is also symmetric, we have $U=V$ and so $U^{\top}U=V^{\top}V=I$ and $UU^{\top}=VV^{\top}=I-N^{-1}{\boldsymbol{1}}{\boldsymbol{1}}^{\top}$. Notice that the mean-field matrix has two parameters, $\phi$ and $\iota$, but, since the leading eigenvalue $\phi^{*}$ is typically easy to be estimated with high accuracy, we may think that the only parameter that is object of inference in this case is $\iota$, which is the quantity that rules the intensity of the “interaction” among the different components due to matrix $\Phi$. We thus obtain $\phi^{*}{\mathbf{v}}^{\top}\Sigma_{det}{\mathbf{v}}=\phi N^{-2}({\boldsymbol{1}}^{\top}\Sigma_{det}{\boldsymbol{1}})$ and, for $2\iota>1$,
${\mathcal{M}}_{det}^{11}=U{\mathcal{S}}_{det}^{11}U^{\top}=\frac{\phi(1-\iota)^{2}}{2\iota-1}UV^{\top}\Sigma_{det}VU^{\top}=\frac{\phi(1-\iota)^{2}}{2\iota-1}(I-N^{-1}{\boldsymbol{1}}{\boldsymbol{1}}^{\top})\Sigma_{det}(I-N^{-1}{\boldsymbol{1}}{\boldsymbol{1}}^{\top}),$
${\mathcal{M}}_{det}^{33}=U{\mathcal{S}}_{det}^{33}U^{\top}=\frac{1}{\phi(2\iota-1)}UV^{\top}\Sigma_{det}VU^{\top}=\frac{1}{\phi(2\iota-1)}(I-N^{-1}{\boldsymbol{1}}{\boldsymbol{1}}^{\top})\Sigma_{det}(I-N^{-1}{\boldsymbol{1}}{\boldsymbol{1}}^{\top}),$
Therefore, recalling $\widetilde{A}_{t}=\tfrac{{\boldsymbol{1}}^{\top}{\mathbf{A}}_{t}}{N}$, $\widetilde{B}_{t}=\tfrac{{\boldsymbol{1}}^{\top}{\mathbf{B}}_{t}}{N}$ and defining $c_{N}=\tfrac{\left({\boldsymbol{1}}^{\top}\Sigma_{det}{\boldsymbol{1}}\right)}{N^{2}}$, the confidence interval based on ${\mathbf{A}}_{t}$ and ${\mathbf{B}}_{t}$ are respectively
$CI_{1-\alpha}(\widetilde{A}_{\infty})\ =\ \widetilde{A}_{t}\ \pm\ \frac{z_{\alpha}}{t^{\phi/2}}\!\sqrt{\phi g(\widetilde{A}_{t})c_{N}}\qquad\mbox{and}\qquad CI_{1-\alpha}(\widetilde{A}_{\infty})\ =\ \phi\widetilde{B}_{t}\ \pm\ \frac{z_{\alpha}}{t^{\phi/2}}\!\sqrt{\phi g(\phi\widetilde{B}_{t})c_{N}}.$
Analogously, the test statistics based on ${\mathbf{A}}_{t}$ is
${{\mathbf{T}}}_{t}=t^{\phi/2}L^{11}(O^{11})^{\top}UV^{\top}\,\frac{{{\mathbf{A}}}_{t}}{g(\widetilde{A}_{t})^{1/2}}=t^{\phi/2}L^{11}(O^{11})^{\top}(I-N^{-1}{\boldsymbol{1}}{\boldsymbol{1}}^{\top})\,\frac{{{\mathbf{A}}}_{t}}{g(\widetilde{A}_{t})^{1/2}},$
and the one based on ${\mathbf{B}}_{t}$ is
${{\mathbf{T}}}_{t}=t^{\phi/2}L^{33}(O^{33})^{\top}UV^{\top}\,\frac{{{\mathbf{B}}}_{t}}{g(\phi\widetilde{B}_{t})^{1/2}}=t^{\phi/2}L^{33}(O^{33})^{\top}(I-N^{-1}{\boldsymbol{1}}{\boldsymbol{1}}^{\top})\,\frac{{{\mathbf{B}}}_{t}}{g(\phi\widetilde{B}_{t})^{1/2}}.$
Remark C.2.
In the particular case $\Sigma_{det}=I$ we have ${\boldsymbol{1}}^{\top}\Sigma_{det}{\boldsymbol{1}}=N$ and hence the following confidence interval based on ${\mathbf{B}}_{t}$:
(C.3) $CI_{1-\alpha}(\widetilde{A}_{\infty})\ =\ \phi\widetilde{B}_{t}\ \pm\ \frac{z_{\alpha}}{t^{\phi/2}}\!\sqrt{\phi g(\phi\widetilde{B}_{t})c_{N}}\qquad\mbox{with}\qquad c_{N}=\frac{1}{N}.$
Moreover, we have ${\mathcal{M}}_{det}^{33}=U{\mathcal{S}}_{det}^{33}U^{\top},$ with
$UU^{\top}=(I-N^{-1}{\boldsymbol{1}}{\boldsymbol{1}}^{\top})\qquad\mbox{and}\qquad{\mathcal{S}}_{det}^{33}=\frac{1}{\phi(2\iota-1)}I,$
which leads to $L^{33}(O^{33})^{\top}=\sqrt{\phi(2\iota-1)}U^{\top}$ and so the test statistics based on ${\mathbf{B}}_{t}$ becomes
${{\mathbf{T}}}_{t}=t^{\phi/2}\sqrt{\phi(2\iota-1)}U^{\top}\,\frac{{{\mathbf{B}}}_{t}}{g(\phi\widetilde{B}_{t})^{1/2}},$
from which we can get
$\|{\mathbf{T}}_{t}\|^{2}=t^{\phi}(2\iota-1)\frac{{{\mathbf{B}}}_{t}^{\top}(I-N^{-1}{\boldsymbol{1}}{\boldsymbol{1}}^{\top}){{\mathbf{B}}}_{t}}{g(\phi\widetilde{B}_{t})/\phi}$ $=\ t^{\phi}(2\iota-1)\frac{\|{\mathbf{B}}_{t}\|^{2}-\frac{({\boldsymbol{1}}^{\top}{\mathbf{B}}_{t})^{2}}{N}}{g(\widetilde{B}_{t})/\phi}$
$=\ t^{\phi}(2\iota-1)\frac{\big{\|}{\mathbf{B}}_{t}-\left(\frac{{\boldsymbol{1}}^{\top}{\mathbf{B}}_{t}}{N}\right){\boldsymbol{1}}\big{\|}^{2}}{g(\phi\widetilde{B}_{t})/\phi},$
that is
(C.4) $\|T_{t}\|^{2}=t^{\phi}2\Delta\frac{\big{\|}{\mathbf{B}}_{t}-\widetilde{B}_{t}{\boldsymbol{1}}\big{\|}^{2}}{g(\phi\widetilde{B}_{t})/\phi}\stackrel{{\scriptstyle d}}{{\longrightarrow}}\chi^{2}(N-1)\qquad\mbox{with}\qquad\Delta=\frac{1}{2}-\frac{\phi^{*}_{2}}{\phi}=\iota-\frac{1}{2}.$
Appendix D Technical results
For the readear’s convenience, we here recall some results used in the above proofs. (For a synthetic review about the notion of stable convergence and its variants, see for instance the appendix of Aletti et al. (2020) and the references recalled therein.)
Theorem D.1.
(A consequence of (Crimaldi, 2009, Theorem A.1))
On $(\Omega,\mathcal{A},P)$, let $(\mathcal{F}_{t})_{t}$ be a filtration and, for each $t\geq 1$, let $(L_{t,k})_{k\in{\mathbb{N}}}$ be a real martingale with respect to $({\mathcal{F}}_{t+k-1})_{k}$, with $L_{t,0}=0$, which converges in mean to a random variable $L_{t,\infty}$. Set
$Y_{t,n}=L_{t,n}-L_{t,n-1}\quad\hbox{for }n\geq 1,\quad U_{t}=\textstyle\sum_{n\geq 1}Y_{t,n}^{2},\quad Y_{t}^{*}=\textstyle\sup_{n\geq 1}\;|Y_{t,n}|.$
[…]
Then, with respect to $(\mathcal{F}_{t})_{t}$, the sequence $(L_{t,\infty})_{t}$ converges to the Gaussian kernel ${\mathcal{N}}(0,U)$ in the sense of the almost sure conditional convergence (and so also stably).
Lemma D.2.
((Aletti et al., 2024, Supplementary material))
If $a_{t}\geq 0$, $a_{t}\leq 1$ for $t$ large enough, $\sum_{t}a_{t}=+\infty$, $\delta_{t}\geq 0$, $\sum_{t}\delta_{t}<+\infty$, $b>0$, $y_{t}\geq 0$ and $y_{t+1}\leq(1-a_{t})^{b}y_{t}+\delta_{t}$, then $\lim_{t\to+\infty}y_{t}=0$.
Theorem D.3.
(A consequence of (Crimaldi and Pratelli, 2005, Proposition 3.1))
Let $({\mathbf{T}}_{t,k+1})_{t\geq m_{0},m_{0}\leq k\leq t}$ be a triangular array of $d$-dimensional real random vectors, such that, for each fixed $t$, the finite sequence $({\mathbf{T}}_{t,k+1})_{m_{0}\leq k\leq t}$ is a martingale difference array with respect to a given filtration $({\mathcal{F}}_{k+1})_{k\geq m_{0}}$. Moreover, let $(s_{t})_{t}$ be a sequence of positive real numbers and assume that the following conditions hold:
  • (1)
[…]
  • (2)
    $\sup_{m_{0}\leq k\leq t}|s_{t}{\mathbf{T}}_{t,k+1}|=s_{t}\sup_{m_{0}\leq k\leq t}\sum_{i=1}^{d}|T_{t,k+1,i}|=O(y_{t})$ with $(y_{t})$ a deterministic sequence such that $\lim_{t\to+\infty}y_{t}=0$.
Then $(s_{t}\sum_{k=m_{0}}^{t}{\mathbf{T}}_{t,k+1})_{t}$ converges stably to the Gaussian kernel ${\mathcal{N}}({\boldsymbol{0}},{\mathcal{M}}_{\infty})$.
Recall that, according to the notation used here, the above condition 2) means that $\sup_{m_{0}\leq k\leq t}|s_{t}{\mathbf{T}}_{t,k+1}|=s_{t}\sup_{m_{0}\leq k\leq t}\sum_{i=1}^{d}|T_{t,k+1,i}|\leq Cy_{t}$ for a suitable constant $C$ and $t$ large enough, so that we trivially have the convergence in $L^{1}$ of $\sup_{m_{0}\leq k\leq t}|s_{t}{\mathbf{T}}_{t,k+1}|$ toward zero, as requested in (Crimaldi and Pratelli, 2005, Proposition 3.1).
The following result combines together stable convergence and almost sure conditional convergence.
Lemma D.4.
(A consequence of (Berti et al., 2011, Lemma 1))
Suppose that $({\mathbf{Y}}^{(1)}_{t})$ and $({\mathbf{Y}}^{(2)}_{t})$ are real random vectors such that the first is adapted to a filtration $({\mathcal{F}}_{t})_{t}$ and the second is measurable with respect to $\mathcal{F}$$_{\infty}$. If $({\mathbf{Y}}^{(1)}_{t})$ converges stably to $\mathcal{N}(0,{\mathcal{M}}^{(1)}_{\infty})$ and $({\mathbf{Y}}^{(2)}_{t})$ converges to $\mathcal{N}(0,{\mathcal{M}}^{(2)}_{\infty})$ in the sense of the almost sure conditional convergence with respect to $({\mathcal{F}}_{t})_{t}$, then
$({\mathbf{Y}}^{(1)}_{t}+{\mathbf{Y}}^{(2)}_{t})\stackrel{{\scriptstyle stably}}{{\longrightarrow}}\mathcal{N}({\boldsymbol{0}},{\mathcal{M}}^{(1)}_{\infty}+{\mathcal{M}}^{(2)}_{\infty}).$